### v2

# Science Prospects for the Southern Wide-field Gamma-ray Observatory: SWGO

Abstract
Ground-based gamma-ray astronomy is now firmly established as a key observational approach to address critical topics at the frontiers of astroparticle physics and high-energy astrophysics. Whilst the field of TeV astronomy was once dominated by arrays of atmospheric Cherenkov Telescopes, ground-level particle detection has now been demonstrated to be an equally viable and strongly complementary approach. Ground-level particle detection provides continuous monitoring of the overhead sky, critical for the mapping of extended structures and capturing transient phenomena. As demonstrated by HAWC and LHAASO, the technique provides the best available sensitivity above a few tens of TeV, and for the first time access to the PeV energy range. Despite the success of this approach, there is so far no major ground-level particle-based observatory with access to the Southern sky. H.E.S.S., located in Namibia, is the only major gamma-ray instrument in the Southern Hemisphere, and has shown the extraordinary richness of the inner galaxy in the TeV band, but is limited both in terms of field of view and energy reach.
SWGO, the Southern Wide-field Gamma-ray Observatory, is an international effort to construct the first wide-field instrument in the south with deep sensitivity from 100s of GeV into the PeV domain. The project is now close to the end of its development phase and planning for construction of the array in Chile has begun. Here we describe the baseline design, expected sensitivity and resolution, and describe in detail the main scientific topics that will be addressed by this new facility and its initial phase SWGO-A.
We show that SWGO will have a transformational impact on a wide range of topics from cosmic-ray acceleration and transport to the nature of dark matter. SWGO represents a key piece of infrastructure for multi-messenger astronomy in the next decade, with strong scientific synergies with the nearby CTA Observatory.
[…]
Contents
- 1 Introduction
- 2 The SWGO Project
- 3 Baseline Design
- 4 Performance
- 5 Galactic Particle Acceleration and Transport
- 6 Transients and Variable Sources
- 7 Particle Physics and Beyond the Standard Model
- 8 Cosmic-Ray Measurements
- 9 Multi-Messenger and Multi-Wavelength Program
- 10 Outreach & Communications
- 11 Summary
1 Introduction
The field of ground-based gamma-ray astronomy has grown rapidly since it was established in 1989 (Weekes et al., 1989). Addressing themes in particle/astroparticle physics as well as high-energy astrophysics, astronomy with photon energies above about eV has emerged as a powerful tool within multi-messenger and multi-wavelength (MM/MWL) astrophysics.
Two complementary ground-based approaches exist: imaging atmospheric Cherenkov Telescope (IACT) arrays, and ground-level particle detectors. The IACTs dominated the early development of the field, but ground-level particle detectors are now firmly established in the Northern Hemisphere. IACTs provide precision and access to the lowest energies, with excellent background rejection power at around TeV energies. However, they are pointed instruments with a modest field of view (FoV) that operate only during darkness. Ground-level particle detection offers a view of the whole overhead sky all of the time, albeit with more modest resolution and without access to the very lowest energies.
The pioneering Milagro detector first demonstrated the potential of the ground-particle approach with the first source detections (Sinnis, 2013). It was clear though that a higher altitude detector was needed to bring the measurement level closer the maximum development depth of air-showers of the relevant energies. In 2015, Milagro’s successor HAWC was completed at 4.1 km altitude in Mexico (Abeysekara et al., 2023). The sensitivity of ground-particle detectors reach a critical level in the HAWC era, where the complementarity of these instruments to the more established IACTs became very clear. According to TeVCat (Wakely & Horan, 2008), HAWC discovered 45 new sources, and added significantly to our knowledge of many more objects. Highlights of HAWC science include constraining the transport of cosmic rays (CRs) in the interstellar medium using nearby pulsars (Abeysekara, A.U. et al. [HAWC Collaboration], 2017a), and establishing jetted microquasars as accelerators up to ultra-high energies (UHE, Abeysekara, A.U. et al. [HAWC Collaboration], 2018b).
In 2021 the LHAASO facility was completed in Daocheng County, Sichuan Province, China, at 4.4 km above sea level. LHAASO contains two ground-level particle detection elements, the Water Cherenkov Detector Array (WCDA) which is a dense array 3.5 times larger than the dense central HAWC array, and the square-kilometer sparse array (KM2A). KM2A consists of buried muon detectors with a 4% fill factor and surface scintillator detectors covering 0.5% of the array footprint. With excellent background rejection power provided by its muon-tagging capabilities and its large area, KM2A has extended gamma-ray astronomy in to the PeV domain (Cao et al., 2021a).
Observations with H.E.S.S., an array of five IACTs in Namibia, have revealed the rich populations of particle accelerators in the inner galaxy (H. E. S. S. Collaboration, 2018b). The Galactic Plane Survey of H.E.S.S. and deep observations of the Galactic Center, do however provide only limited sensitivity beyond 50 TeV. The CTA Observatory (CTAO) is planned to include a southern site in Chile and will dramatically improve the survey depth (Abe et al., 2024) and resolution (see e.g. Schwefer et al., 2024). However, no major wide-field VHE/UHE instrument exists so far in the southern hemisphere, although developments are underway in Bolivia towards ALPACA, a UHE-focused instrument (Anzorena et al., 2024).
A wide-field observatory located in the Southern Hemisphere, with broad energy range and with sensitivity comparable to that of LHAASO would strongly complement CTAO, in particular in the areas of transient phenomena, large scale emission and UHE sources. Here we describe the expected scientific performance of a new ground-level particle detecting gamma-ray observatory, planned for construction in the Southern Hemisphere and based on water Cherenkov detectors (WCDs): SWGO.
2 The SWGO Project
The Southern Wide-field Gamma-ray Observatory (SWGO) Collaboration was created in 2019, as a convergence of several earlier concepts and efforts including LATTES (Assis et al., 2018), and the SGSO Alliance (Albert et al., 2019). The central concept for the observatory was defined as
1.
A gamma-ray observatory based on ground-level particle detection, with close to 100% duty cycle and order steradian FoV.
2.
Located in South America at a latitude of -30 to -10 degrees.
3.
Placed at an altitude of 4.4 km or higher.
4.
Covering an energy range from 100s of GeV up to the PeV scale.
5.
Based primarily on WCD units.
6.
Utilizing a high fill-factor core detector with area considerably larger than HAWC and significantly better sensitivity, alongside a low density outer array.
7.
Modular and scalable, with the possibility of extensions and/or enhancements.
WCD units are attractive in terms of cost per unit area of particle detection, and were the approach of choice for Milagro, HAWC, and two of the three ground-particle components of LHAASO. From the beginning the SWGO Collaboration has been committed to close scientific coordination with the CTA Observatory, recognizing the very strong synergy and complementarity between these instruments.
The SWGO Collaboration has now grown to include over 100 institutes in 17 countries on four continents, and is close to completion of its R&D Phase, with milestones defined at the beginning of the project, reached over the last few years (see Table 1). Notable amongst these milestones was the establishment of a baseline approach based on extensive simulations covering a wide phase space (M6: see Conceição, 2023, and Section 3), as well as design development and cost estimation work. Another critical milestone was the identification of a preferred site for SWGO (M7), based on criteria of scientific performance (including altitude and latitude effects), cost (including construction and a nominal 5 years of operations), and risks. This milestone was reached in July 2024 (see Section 3.1) after a comprehensive process that considered 10 sites in four countries Santander et al. (2023).
| Milestone | Completed | |
| M1 | R&D Phase Plan Established | Q1 2020 |
| M2 | Science Benchmarks Defined | Q2 2020 |
| M3 | Reference Configuration & Options Defined | Q4 2020 |
| M4 | Site Shortlist Complete | Q3 2022 |
| M5 | Candidate Configurations Defined | Q1 2022 |
| M6 | Performance of Candidate Configurations Evaluated | Q3 2023 |
| M7 | Preferred Site Identified | Q2 2024 |
| M8 | Design Finalised | - |
| M9 | Construction & Operation Proposal Complete | - |
The optimization process has been guided by science benchmarks associated with key science cases, established early in the project (M2) and allowing us to objectively optimize the instrument design based on our scientific objectives. The benchmarks cover the topics of transient sources (specifically GRBs, see Section 6.2), Galactic PeV particle accelerators (see Section 5.3), extended sources (with the example of PWNe and associated halos, see Section 5.4.1), large-scale diffuse emission (see Section 5.2.1), dark matter detection (see Section 7.1) and the mass-resolved anisotropy of CRs (see Section 8).
Figure 1 shows the sky visible to SWGO, including the Fermi Bubbles and the whole inner galaxy. The Galactic Centre passes directly overhead at the SWGO Site.
SWGO construction is anticipated to take place in stages. After a modest scale Pathfinder, the first science-capable phase is anticipated to be SWGO-A, which will focus on the low to intermediate energy range (see Section 4).
3 Baseline Design
The SWGO Collaboration is now approaching Milestone 8 - the definition of the Baseline Design. The array site has been chosen, the overall layout has been defined, and the detector unit design is fixed at least for the inner part of the array. In this section we give a brief overview of these aspects of the design, discuss areas where options still exist and note the existence of already funded and possible future extensions to the project.
3.1 The SWGO Site
The SWGO primary site – Pampa La Bola – is located in the Calama commune in northern Chile, at in altitude of 4,770 meters. The site lies on an extensive plateau within the Atacama Astronomical Park (Ocampo et al., 2020), in close proximity to several major astronomical facilities, including the Atacama Large Millimeter/submillimeter Array (ALMA). The site has excellent access, only 45 minutes drive from the town of San Pedro de Atacama, which lies at 2.4 km altitude. Figure 2 illustrates the location of the site within Chile, the local region and the array layout (see Section 3.2), as well as a photograph showing the excellent ground conditions. The site lies close to the international road Route 27.
The site was selected from three shortlisted candidate sites in Peru, Argentina and Chile in 2024, following an extensive search campaign (Santander et al., 2023). The back-up SWGO site is located close to the town of Imata, in the Arequipa district of Peru.
The first activities planned at the site include building the Pathfinder (SWGO-PF), aimed to demonstrate the validity of the chosen technologies and assembly and operation procedures. The SWGO-PF will consist of 4-6 WCDs plus associated control and readout, water storage and purification and power systems. The main purpose of the SWGO-PF is to gain experience and capture lessons learned for the installation of the tanks, for the water transport and quality, to demonstrate system integration at high altitude and to validate our simulations.
The next phase of the project is planned to be SWGO-A, the first component of the high fill factor innermost array zone (160 m, see Figure 2). SWGO-A will be constructed next to the Pathfinder location to benefit from the initial infrastructure.
3.2 SWGO Array Layout & Simulation
The SWGO detector is made up of an array of independent WCD units. These units are distributed in a 3-zone configuration (see Figure 2). The central zone consists of closely packed tanks with only enough separation to accommodate construction and maintenance. The fill factor of the central zone is 70%. This high-density region provides the instrument’s sensitivity from hundreds of GeV to tens of TeV gamma rays, and is fundamental to the sensitivity for transients and detection of distant gamma-ray emitters. Surrounding the central zone are two outer zones of gradually decreasing fill factor: 4% for the intermediate zone and 1.7% for the outer zone. These outer zones extend the total instrumented area to 1 km2 and provide detection capability up to the PeV range. In total the configuration contains 3763 WCDs.
Whilst the inner array WCD design has been baselined, several options are under consideration for the outer array (see Section 3.5).
3.3 WCD Unit Design
The baseline for the central array WCD unit is a cylindrical steel tank of radius 5.2 m and height 4.1 m, assembled on site and with a bladder to contain the water. The bladder has a total height of 4.0 m and comprises two volumes, with the lower section measuring 80 cm in height, and a Tyvek surface layer for high reflectivity (see Kunwar et al. (2023) for a study on the dual-layer approach and the impact of liner reflectivity). A photograph of an assembled prototype 5.2 m tank at the HAWC observatory can be seen in Figure 3, together with a photograph taken inside a deployed unit and a simulated muon passing through the WCD. The lower chamber provides the unit with muon-tagging capability to aid in rejection of the charged cosmic-ray background.
Once filled with filtered water, a double-PMT unit is lowered into a double-chamber bladder to position a downward-looking 8-10" PMT in the lower section and a 10" PMT looking up from the bottom of the upper section. We are currently evaluating two options for high-voltage (HV) supplies to the PMT: active (locally generated HV, similar to the active base used for IceCube) and passive (signal and HV carried on the same cable).
Signals from the PMTs in each WCD are routed to nearby Field Nodes (FNs), which serve as digitization and control units. Each FN nominally serves 55 WCDs, corresponding to 110 PMT channels. Housed in standalone, environmentally protected cabinets, the FNs include heating, cooling, insulation, and a programmable logic controller (PLC) for local control and monitoring. PMT signals are digitized using FlashCam electronics, a system developed for the CTAO (Werner et al., 2017), which operates at 250 MSa/s with deep waveform buffers. Digitization takes place following a first-level trigger, based on the signal from any PMT exceeding a fraction of a photoelectron, thereby requiring a low level of electronic noise.
Depending on the HV configuration, signal routing differs. In the case of passive HV distribution, a single coaxial cable carries both the PMT signal and HV from the WCD to the FN. Inside the FN, dedicated PhantomHV boards (Haist, 2023) generate the HV and separates it from the signal path. The resulting signal is routed to the FlashCam digitizers. The PhantomHV boards include a Time-Domain Reflectometry. system to measure the signal propagation delay along the coaxial cable to each PMT, allowing for per-channel cable transit time correction. This is used to compensate for temperature-dependent variations in cable delay and can also be used to detect faults or damage in the cable. The system achieves sub-nanosecond precision, supporting accurate event timing across the array, as required for precise direction reconstruction.
In configurations with locally generated HV at the WCD, Cat6 cables carry the PMT signal, low-voltage power, and slow-control communication directly between the WCD and FN. In this case, electronics within the FN split the signal and auxiliary lines, with the signal again routed to FlashCam for digitization.
Each FN includes a redundant White Rabbit (WR) timing system, which provides sub-nanosecond synchronization across the array. WR combines IEEE 1588 Precision Time Protocol with Synchronous Ethernet and phase-tracking techniques to distribute a common time and frequency base over dedicated fiber links. Each FN is connected to a central location via fiber for timing, control and raw data readout. Power is supplied via a single-phase AC feed.
Centrally, the approximately 50 FNs needed for the central array, are connected to a computing rack for control and data acquisition. Several WR switches, with one acting as a master node, are connected to a GPS receiver to provide time synchronization across the array. An array-level trigger is formed in software using the digitized waveforms in real time.
All elements of the inner array baseline have been prototyped, and the simulated performance presented below is based on this design. For the outer array, alternative solutions are under exploration, see Section 3.5 below.
3.4 The first stage of SWGO: SWGO-A
The SWGO-A layout corresponds to a 385 WCD-unit segment of the inner array, with a fill factor of 65%, arranged in seven independent clusters of 55 tanks. Each cluster will be electrically connected to a single FN as seen in Figure 4. The WCDs will be separated by 50 cm, permitting access to all sides. Between every two rows of tanks, a 1.5 m wide aisle is planned, for installation of the field node, access by small cranes, or similar equipment as needed for deployment and maintenance. The overall fill factor is slightly larger than that of the central array of the HAWC observatory.
3.5 Outer Array Options
The outer array plays a vital role in increasing the effective detection area, enhancing background rejection, and boosting the overall performance of the observatory at higher energies. Due to the large footprint of this element of the array, alternative tank configurations are currently being explored. The final design is still under development and is undergoing an optimization process aimed at improving hadronic background discrimination, aiding the reconstruction of shower cores for events falling outside the central array, and increasing sensitivity to high-energy gamma rays in the multi-TeV to PeV range. The WCD design used in the central array is also an option for the outer array. However, the solution must also be cost-effective for covering large areas and therefore may employ simpler or smaller tanks than those used in the central array.
In the case in which the outer array is based on the WCD design presented above for the central array, the size and location of field nodes, as well as cabling, will need to be optimized.
One promising alternative involves the use of smaller High Density Polyethylene (HDPE) roto-molded plastic tanks. Prototype tanks, based on a redesigned version of the tanks used at the Pierre Auger Observatory, have been manufactured by Rotoplastic. Whilst studies so far indicate an advantage for deeper tanks in terms of background rejection power, smaller tanks may be cost effective in the outer array and are particularly interesting at the SWGO site given the significant cost of transporting water. Such tanks, measuring 3.6 meters in diameter, could be produced near the observatory site and transported by standard trucks.
Compared to the original Pierre Auger design, several improvements are being considered: the use of enhanced resin materials for greater mechanical strength and UV resistance under harsh Atacama conditions; a potential double-wall structure for thermal insulation to prevent water freezing; and the inclusion of a new Reflective Oxide Nanocoating as the inner layer of the tank. This could eliminate the need for a Tyvek-laminated polypropylene bladder, resulting in a WCD unit that may be more cost-effective, durable, and easier to deploy at high altitudes.
Some full-scale prototypes have been successfully produced and shipped to Chile for SWGO-PF (see Figure 5), including one with a reflective nano-coating and one that is in operation equipped with a multi-PMT module (Figure 6). This module is composed of 73” PMTs to maximize the angular acceptance of Cherenkov light produced by air-shower particles that penetrate the tank, and to provide directionality information at each WCD unit.
Currently, three main configurations are being considered:
1.
A sparse array of WCDs identical to the ones of the central array but spaced further apart (tens of meters).
2.
A sparse array of single-layer, smaller HDPE WCD tanks, with a water depth of 1.7 m, potentially instrumented with a multi-PMT module.
3.
A sparse array of double-layer, smaller HDPE rotomolded plastic tanks. The exact dimensions of the two layers and the type and arrangement of the photosensors are under study.
The outer array will make use of a WR time synchronization system common to the one used for the inner array.
3.6 The Radio Array SWGO-TURBO
An array of 800 radio antennas, named The Utility for Radio Beam-formed Observations (TURBO), will be constructed as part of the SWGO Project, following a successful bid to the ERC Consolidator Grants. The dual-polarized antennas will observe the pulsed radio emission from air showers in the 50-200 MHz frequency band, and data acquisition will be fully integrated within the observatory. Since SWGO-TURBO is already funded, it will be deployed during the early phase(s) of SWGO and therefore complement the particle observations for the full operational time of the observatory. The readout of the radio array will be initiated by a trigger from the particle detectors. By applying interferometric reconstruction Schoorlemmer & Carvalho (2021); Schoorlemmer, H [Pierre Auger Collaboration] (2023), it is anticipated that the influence of noise can be significantly reduced compared to traditional radio observations of air showers.
The primary focus of TURBO is on air showers with high energy (PeV) and large zenith angles (between and ). For these air showers, radio will complement the array of WCDs by providing accurate angular reconstruction, an alternative energy estimation, and reconstruction of the depth of the air shower maximum. These observations aim to increase the sensitivity and aperture of SWGO for photons with energy above PeV and enable more detailed air shower observation of charged CRs.
3.7 Possible UHE Extension in a Lake
The deployment of WCDs inside a lake was considered as part of the SWGO design study ( Goksu et al. (2025); Goksu & Hofmann (2021)), given that lake water provides good shielding of electromagnetic particles, with the potential to enhance muon tagging capabilities.
No suitable high-altitude lake was found, but investigations continue within the collaboration, targeting a somewhat lower altitude and a low fill factor array aimed at ultra-high energies. As well as searching for candidate lakes, the collaboration is developing lake deployment technologies and will compare the cost-effectiveness of a lake extension at a second site in comparison to expansion of the primary site.
4 Performance
This section, discusses the expected performance of SWGO using the reference layout described in Section 3. The areas of the three SWGO zones with fill factors of 70%, 4%, and 1.7%, decreasing from the center, are shown in Figure 2. Figure 4 shows the layout of SWGO-A, which has a fill factor of 65%. The same WCD unit design was simulated for the three zones, and also for SWGO-A. The performance of SWGO at an altitude of 4770 m is examined using simulations based on CORSIKA (Heck et al., 1998) and HAWCsim (Abeysekara, A.U. et al. [HAWC Collaboration], 2017b). Air showers of secondary particles induced by gamma rays and protons are simulated with energies from 100 GeV to 1 PeV and utilized to derive Instrument Response Functions (IRFs) using the pyswgo software framework. The event reconstruction is partly carried out with AERIE (Abeysekara, A.U. et al. [HAWC Collaboration], 2018a, 2017b), originally developed for HAWC.
To ensure precise observations, only events with a multiplicity of at least 65 stations are accepted for SWGO111By the same standard, 30 triggered stations are required for SWGO-A., which corresponds to a trigger threshold at a signal-to-noise ratio similar to that of HAWC. Furthermore, events with reconstruction uncertainty that deviate by more than two standard deviations from the expected energy reconstruction or the angular reconstruction resolution are rejected.
4.1 Limitations
The current simulation setup does not consider the background from air showers (coinciding/overlapping showers, single muons) nor from dark counts. The impact of these backgrounds on the analysis can be considerably reduced with dedicated filtering methods, but for the moment, they could limit performance at energies close to the trigger threshold. At higher energies, the current IRFs are limited by the strong rejection power of the gamma-hadron separation, causing the sensitivity estimates to feature a significant statistical uncertainty due to very limited or vanishing Monte Carlo samples. This leads to conservative estimates on the performance at energies larger than 100 TeV, which applies to all science benchmarks discussed in this document. As the project continues to consider the layout and unit design of the outer array, adjustments in performance are anticipated. Furthermore, we use state-of-the-art machine learning algorithms (Erdmann et al., 2021) during gamma-hadron separation. Although their performance on simulations is exceptional and has been validated using measured data in CR physics (A. Abdul Halim et al.(2025b), Pierre Auger Collaboration; A. Abdul Halim et al.(2025a), Pierre Auger Collaboration) and neutrino astronomy (R. Abbasi et al.(2023), IceCube Collaboration), a validation of these novel algorithms in gamma-ray astronomy is still lacking, and will be crucial for achieving the performance shown here.
The derived quantities discussed hereafter provide an approximate performance estimate of SWGO and SWGO-A based on detailed simulations, dedicated reconstruction software, and reconstruction algorithms tailored to the SWGO design. Statistical uncertainties are shown as opaque regions and are obtained by bootstrapping.
4.2 Gamma-hadron separation
Gamma-hadron separation is performed by training a shallow neural network given several separation observables in bins of shower energy, zenith, and core distances with respect to the array center, accounting for the graded design of SWGO. In addition to low-level observables (e.g. zenith angle and number of triggered stations, and high-level observables such as PINCness and LDFChi2 introduced by HAWC, R. Alfaro [HAWC Collaboration] et al., 2022), the analysis also makes use of novel deep-learning-based algorithms, such as graph neural networks (Glombitza et al., 2025a) and transformers (I. Watson, HAWC Collaboration, 2023), and uses the output scores of those algorithms in classification. The performance of the gamma-hadron separation is shown in Figure 7 for events with zenith angle between and for SWGO (orange) and SWGO-A (grey), and compared to HAWC (green) (A. Albert et al., 2024) and LHAASO222At the time of writing, no performance data is available for LHAASO WCDA KM2A (blue) (LHAASO Collaboration, 2024b). Dashed, dotted, and dash-dotted lines indicate different zones of the SWGO layout. In the top panel, the -efficiency (true positive rate) is shown and has been set to of the triggered events. Additional selection cuts slightly decrease these efficiencies. The background rejection (false positive rate), shown in the bottom panel, strongly decreases as a function of energy and reaches a suppression of and lower for all zones, demonstrating the efficacy of the double-layer approach. SWGO-A reaches a rejection of at low energies, and approaches above 10 TeV, showing significant improvements compared to HAWC . Compared to LHAASO KM2A, SWGO exhibits superior background rejection performance until 100 TeV at which — due to the strong rejection and the limited amount of Monte Carlo statistics — no background remains in any of the investigated zones and layouts.
In Figure 8, we show the effective area for detecting gamma rays after gamma-hadron separation using SWGO (orange) and SWGO-A (grey) as a function of energy for events with zenith angle between 0∘ and 30∘. As a comparison, we also show HAWC (green) (A. Albert et al., 2024), LHAASO WCDA (light blue) (LHAASO Collaboration, 2021), and LHAASO KM2A (blue) (LHAASO Collaboration, 2024b). The thin orange lines denote the effective areas of events reconstructed with the shower core in the three different SWGO zones. Whereas SWGO zone 0 and SWGO-A become fully efficient at 1 TeV, zones 1 and 2 are efficient above 10 TeV and 30 TeV, respectively. The continuous orange line denotes the total effective area of the entire detector. With its m2 size, SWGO’s graded-zone design enables efficient gamma-ray detections from 300 GeV to the PeV scale, surpassing LHAASO KM2A. Similar to HAWC, we included off-array events in this work to increase the effective area of SWGO-A, reaching almost m, close to LHAASO WCDA and HAWC. Due to its high-altitude location, already SWGO-A lowers the threshold compared to HAWC and LHAASO WCDA, providing promising potential for improved observations at low energies.
4.3 Event reconstruction
The utilized angular reconstruction is based on the Gaussian fitting procedure introduced by HAWC (Abeysekara, A.U. et al. [HAWC Collaboration], 2017b), in which a plane with a second-order curvature model of a shower front is fit to arrival times. The angular resolution, defined as the quantile of the angular distance, is shown in Figure 9 left for SWGO-A (grey) and SWGO (orange) and compared to various operating ground-based gamma-ray observatories. While SWGO-A shows a resolution very similar to HAWC (green) (A. Albert et al., 2024) and LHAASO WCDA (light blue) (LHAASO Collaboration, 2021), SWGO demonstrates a very precise reconstruction with significant improvements over LHAASO and HAWC. After running the angular fit, the shower core and energy reconstruction are carried out using template-based methods (Parsons & Hinton, 2014; Joshi et al., 2019) adapted to the SWGO design (F. Leitl, SWGO Collaboration, 2023), with the arrival direction serving as seeding parameters. These state-of-the-art template methods utilize likelihood models of the signals in tanks as a function of core distance in different bins of energy and arrival direction, which are derived using a comprehensive simulation library. By maximizing the likelihood of the observed event, template methods offer reliable performance across a large phase space, thereby significantly improving the reconstruction compared to previous approaches. The quality of the energy reconstruction is summarized in Figure 9 right, for SWGO-A (grey) and the three different zones of the SWGO array (orange). The energy resolution, defined as the quantile of the difference in log space, is improving significantly as a function of energy. Above 1 TeV SWGO (SWGO-A) reaches a resolution better than 0.2 (0.3) for all zones, improving to 0.05 (0.1) at the highest energies. The core resolution, defined as the 1D 68% quantile of the distribution of the distance between the reconstructed and the true core position, improves from about 25 m to around 2 m and is below 10 m for energies larger than 3 TeV, providing accurate event reconstruction binned in different zones.
Based on the derived IRFs and the Gammapy software (Donath et al., 2023), Figure 10 shows the one-year (dashed) and five-year (continuous) differential sensitivity for SWGO (orange) and SWGO-A (grey) to detect a steady point source with an energy spectrum of with significance at the level of for a continuously emitting source at declination. Whereas lines represent the expected performance estimated using the simulated IRFs, the dotted regions indicate extrapolations of the sensitivity in the phase space where no background events pass the gamma-hadron separation, assuming vanishing signals and constant rejection performance. Note that these extrapolations are currently not considered in the IRFs used for the science benchmark, limiting their expressiveness beyond 100 TeV. For comparison, show the sensitivity of LHAASO (blue) (Vernetto, S for the LHAASO Collaboration, 2016), HAWC (green) (A. Albert et al., 2024), with one-year observation time, as well as 50 hours of observation time using CTAO (purple) (The CTA Consortium, 2019). Utilizing its innovative detector design at high altitudes, SWGO-A provides distinct improvements at low energies compared to HAWC. The expected SWGO sensitivity shows very promising performance over the entire energy range, with prospects for significant improvements over LHAASO at low and medium energies, demonstrating its strong potential to complement CTAO in the TeV range. As the layout and unit design of the outer array will be considered in the future, improvements are expected at the highest energies. With the shown sensitivity, the SWGO design will enable us to survey the Southern sky with unprecedented precision.
5 Galactic Particle Acceleration and Transport
5.1 Unbiased Survey of the Galactic Plane
Previous generations of instruments have detected several populations of Galactic sources including pulsar wind nebulae (PWNe), supernova remnants (SNR), star-forming regions (SFRs), TeV halos, pulsars, gamma-ray binaries and micro-quasars (see Section 5.4 for dedicated discussions). A large fraction of new detections arose from surveys by HEGRA, H.E.S.S., VERITAS, Milagro, HAWC, LHAASO, and the Fermi Gamma-ray Space Telescope, highlighting the importance of unbiased observational strategies. By providing large and unbiased samples of gamma-ray sources, surveys help to uncover the mechanisms of particle acceleration, their relative efficiencies, the nature of their transport, and the contributions of different source populations to the galactic CR population.
The SWGO design is exceptionally well-suited for unbiased surveys of the Galactic Plane due to its wide FoV and high duty cycle. It will allow a large fraction of the sky to be continuously monitored, including most of the inner galaxy where the majority of galactic sources are concentrated.
To test the survey capabilities of SWGO, 1 to 10 years of Galactic Plane observations were simulated between Galactic longitude and latitudes . The source population and interstellar emission models used are the same as for the recent CTA Galactic Plane Survey (CTA-GPS) simulation333These models are publicly available at https://zenodo.org/records/10008527 (Abe et al., 2024). This includes both a compilation of known sources detected at GeV or TeV energies and synthetic populations for the main classes of galactic VHE sources (SNRs, PWNe, and gamma-ray binaries). Section 5.2.1 discusses the specifics of the alternative interstellar emission models considered and the capabilities of SWGO to detect diffuse emission. A potential view of the Galaxy above 1 TeV as seen by SWGO after 5 years of observation is given by Figure 11, which shows the significance of the excess above the CR background integrated in a 0.1∘ correlation radius.
The number of expected detections above a given test statistic (TS) value, is computed as a benchmark quantity. For simplicity this number is estimated only for the Asimov data set (i.e counts equals predicted counts, see Cowan et al., 2011), which gives a limit on the expected detection for the true simulated model. For this benchmark a minimal interstellar emission model (IEM) is used as in the CTA-GPS simulation, which yields an upper estimate for the number of detectable sources.
In Table 2, the expected detections from SWGO are compared to the ones expected for the CTA-GPS, and to the catalogs from the current generation of instruments such as H.E.S.S., HAWC, and LHAASO. This result clearly shows that SWGO will have a transformational impact on the field by multiplying the number of detected sources by up to a factor of 3 in a single year and up to a factor of 6 in 10 years.
In this paper there is not attempt to derive a catalog from scratch as was done for the CTA-GPS simulation analysis (see Appendix D of Abe et al., 2024), but the same techniques could be applied. Above 10 TeV where the sensitivities and angular resolution are comparable, similar performances could be expected. At lower energies, where both spectral and angular resolution are degraded, larger confusion effects are expected so it will be more difficult to identify individual sources, or to study energy-dependent morphology. But these issues could be alleviated by performing joint analyses with CTAO.
Table 2 also demonstrates that the expected number of detections of the CTA-GPS is matched by SWGO in about 5 years, while so far the scheduling of the CTA-GPS is proposed to be spread over 10 years. The presence of SWGO may therefore have a direct impact on CTAO observation strategy. Regarding the synergy between SWGO and CTAO, among the sources detectable by whether SWGO or CTAO, about 70% would be significantly seen by both instruments independently. As SWGO and CTAO are both committed to deliver open data this would offer great opportunities for joint analyses to the community, and even the possibility to build a joint catalog of gamma-ray sources over four decades in energy. This could even be extended to six decades in energy, with the inclusion of Fermi Large Area Telescope (LAT) data.
| Catalogue | Detected sources |
| HGPS(a) | 96 |
| 3HWC(b) | 48 |
| 1LHAASO(c) | 54 |
| Future surveys | Expected detections |
| CTA-GPS(d) | 461 |
| SWGO-A 1 yr | 70 |
| SWGO-A 5 yrs | 135 |
| SWGO 1 yr | 359 |
| SWGO 5 yrs | 487 |
| SWGO 10 yrs | 536 |
| SWGO 10 yrs + CTA-GPS(e) | 603 |
5.2 Extended Galactic Structures
The study of astrophysical sources with extended VHE emission has become a central topic in modern astrophysics. In particular, morphological characterization provides critical insights into the physical mechanisms driving particle acceleration and propagation. SWGO is designed as a survey instrument with a large FoV, to efficiently map substantial portions of the Galactic disk and to study extended sources. Its high sensitivity at VHE, along with an enhanced spatial, and energy resolution will enable the investigation of extended sources and the diffuse emission of the Galactic Plane.
5.2.1 Galactic Diffuse Emission
The local measured CRs up to the Knee or even up to EeV are believed to originate from Galactic sources. However, information about CR sources and the CR energy distribution at the source is lost as soon as CRs start to diffuse in the interstellar medium (ISM). While escaping the accelerating source and diffusing in the Galaxy, CR protons and nuclei interact inelastically with the atoms and molecules of the interstellar gas and produce gamma rays through the decay of neutral pions. Electrons emit gamma rays through inverse Compton (IC) scattering off the radiation fields and through bremsstrahlung (Hunter et al., 1997; Moskalenko et al., 1998; Strong et al., 2000; Aharonian & Atoyan, 2000; Kelner et al., 2006; Dundovic et al., 2021; Porter et al., 2022). In contrast to CRs, the gamma rays produced by these accelerated particles are neutral and thus travel unperturbed from the emission site to the detector. The gamma-ray energy flux from different directions on the sky thus provides direct information about the parent CR flux and the CR transport properties in different locations in the Galaxy.
Although multiple studies of the diffuse emission have been performed at GeV energies using multiple missions and recently Fermi-LAT (Hunter et al., 1997; Ackermann et al., 2012; Acero et al., 2016), measurements at TeV energies are more challenging because the CR spectrum and the resulting diffuse emission spectrum is steeper at higher energies (Amenomori et al., 2021; Cao et al., 2023b; Yan et al., 2024; Alfaro et al., 2024a; LHAASO Collaboration, 2024a). Furthermore, it is difficult to separate the faint large-scale galactic diffuse emission from the residual hadronic background and possible contamination from unresolved sources.
While HAWC (Alfaro et al., 2024a) and LHAASO (Cao et al., 2025a) have deeply surveyed the northern sky, including measurements of the diffuse emission, SWGO will extend the survey to the whole southern sky with improved angular and energy resolution. SWGO’s superior gamma-hadron separation provides better sensitivity than current instruments.
In Figure 12 the diffuse emission measurement by LHAASO is compared to different IEMs, in particular “CRINGE” (Schwefer et al., 2023), and “varmin-rescaled” (Abe et al., 2024) that feature homogeneous and inhomogeneous CR transport, respectively. Interestingly, in this large region, which is far from the contamination of known sources, as considered in previous studies by LHAASO, the two models are barely distinguishable. Moreover, once the contribution from the unresolved source population is added, both models can account for the diffuse emission flux measured by LHAASO within uncertainties. In order to distinguish between these two classes of models, and further study the CR transport properties, it is necessary to probe the innermost region of the Galaxy at smaller scales, which should be achieved with SWGO.
Indeed SWGO will have the capacity to detect diffuse emission down to scales of in the inner galaxy, and in the outer galaxy as shown in Figure 13. It can therefore resolve individual molecular clouds (MCs) in the inner galaxy at TeV energies, like Fermi-LAT at GeV energies. SWGO will detect gamma rays from giant MCs satisfying where is the cloud mass in 10 and is its distance in kpc, assuming that the MC is illuminated by a CR spectrum similar to local measurements (Peron & Aharonian, 2022). Figure 14, shows as an example that SWGO would be capable to detect the Bania clump 2 (Bania, 1977; Stark & Bania, 1986), a famous cloud of the inner galaxy, and probe its spectrum up to 300 TeV. In regions where source contamination is negligible, as in this example, SWGO would not only detect individual MCs but also be able to discriminate between models featuring different CR transport properties.
5.2.2 Fermi Bubbles
The Fermi Bubbles are among the most prominent large-scale gamma-ray structures in our galaxy. These massive lobes extend nearly 50∘ above and below the Galactic Center (GC), spanning a total diameter of approximately 10 kpc (Su et al., 2010; Ackermann et al., 2014a). Their characteristic hardness (around in the gamma-ray range) challenges our understanding of particle acceleration processes occurring within the Milky Way, and suggests the presence of highly energetic particles that interact with interstellar matter and/or radiation fields (Yang et al., 2018; Crocker & Aharonian, 2011; Sarkar, 2024).
Both the origin and the nature of the radiating particles in the Fermi Bubbles are important open questions in high-energy astrophysics. In fact these two questions are closely coupled. The competing scenarios for energizing the bubbles are associated with different timescales and hence different physical processes.
In one scenario, the central supermassive black hole of our galaxy (Sgr A*) went through a period of much greater accretion a few million years ago, leading to AGN-like jet activity that injected a significant amount of energy into the Milky Way’s halo (Su et al., 2010; Yang et al., 2018). This scenario suggests that the Fermi Bubbles are relics of a past AGN phase, with accelerated particles producing gamma rays through IC scattering and hadronic interactions. Observations indicate that the Fermi Bubbles exhibit a surprisingly hard gamma-ray spectrum from 1 GeV to 100 GeV, with no detected spectral cutoff below 100 GeV. This spectrum remains relatively uniform across both northern and southern lobes, which is difficult to explain within standard galactic CR propagation models.
The alternative scenario involves much slower inflation of the bubbles powered by ongoing star formation activity, and associated particle acceleration, over Gyr timescales. In this case, the bubbles are driven by large-scale winds originating from intense nuclear starburst activity at the GC (Su et al., 2010; Yang et al., 2018). The very long radiative timescales for accelerated protons and nuclei in the bubbles are then matched to the overall lifetime of the system and become efficient in terms of gamma-ray production.
SWGO is uniquely suited to explore the VHE and UHE regimes of the Fermi Bubbles, thanks to its continuous sky coverage and exceptional sensitivity to extended gamma-ray emission. Its strong performance at the declination of the Fermi Bubbles offers an unprecedented opportunity to study these structures in greater detail. Such observations could yield new insights into the nature of the Fermi Bubbles, advance our understanding of CR acceleration and propagation, and shed light on the activity history of the supermassive black hole at the center of our galaxy.
5.2.3 The Galactic Center and the Central Molecular Zone
The center-most 200 pc of our galaxy is a very rich region populated by: dense MCs forming the central molecular zone (CMZ), SNRs like the composite SNR G0.9+0.1, massive stellar clusters like the Quintuplet, the Arches, and the Nuclear clusters, where active star formation is ongoing, and the supermassive black hole Sgr A⋆. The complicated image of the GC has been a challenging region for a long time and disentangling its sources and processes will enhance our knowledge of the physics active, not only the center of our galaxy, but potentially all the barred spiral galaxies in the universe.
Abramowski, A. et al. (2016) (H.E.S.S. Collaboration) was the first to report detection of a candidate PeVatron accelerator at the GC (see Section 5.3 for further discussion on this science case). The spectrum of the central source in coincidence with Sgr A⋆, was extracted from a circular region of radius 0.1∘ centered on Sgr A⋆. The best fit spectrum is an exponential cutoff power-law (ECPL) spectrum with a photon index of , and an energy cut-off at TeV. For diffuse emission around the central source, an extended annulus centered at Sgr A⋆ was adopted with radii between 0.15∘ and The best-fit spectrum of this extended region is given by a power-law (PL) with photon index , and an ECPL model is not preferred (per Abramowski, A. et al. (2016), H.E.S.S. Collaboration). Figure 15 demonstrates the respective flux points and spectra of these two sources.
HAWC detected the GC after seven years of observation Albert et al. (2024), and improvements to the algorithms for event reconstruction and gamma-hadron separation methods for high zenith angle observations. Although the HAWC image of the GC has lower angular resolution, the central source HAWC J17462856 can be modeled as point-like, and the best fit spectrum is a PL with photon index , shown in Figure 15. Due to the lower resolution, it is difficult to separate the two point-like sources detected by H.E.S.S. (HESS J1745290 the central source, and HESS J1746285 know as the Arc source) from the diffuse emission. Hence, the flux points in Figure 15 do not consider such a distinction. However, an expected flux can be calculated by subtracting the two sources (see Figure 1 in Albert et al., 2024), which shows a very similar photon index. Note that the photon index of the central source HAWC J17462856 is much steeper than the photon index of the diffuse emission reported by the H.E.S.S. collaboration. This might indicate a cut-off at energies beyond the H.E.S.S. optimal range (tens of GeV to tens of TeV), where it starts to lose statistics.
With its location in the southern hemisphere, SWGO will be the optimal ground particle detector to observe the GC region, which passes right overhead, and will reach energies up to hundreds of TeV. It will detect the GC region in less than one year of observation, as shown in Figure 15, and provide a clearer spectral shape in the highest energy range, especially for the diffuse emission of the CMZ. These advances are crucial to understand the particle acceleration physics in the GC, could also better describe the three dimensional distribution of the dense MCs in the CMZ.
5.3 Ultra-High-Energy Sources and PeVatrons
The concept of a “PeVatron” refers to astrophysical accelerators of particles — either electrons, positrons or protons — to energies of at least 1015 electronvolts (PeV). This definition is fundamentally determined based on the largest rigidity ( with the atomic charge) that particles can attain at a given acceleration site. PeVatrons can be divided into two categories: “hadronic PeVatrons” and “leptonic PeVatrons”. In hadronic PeVatrons, gamma-ray emission predominantly results from proton-proton (pp) and proton-photon (p) interactions via the decay of neutral pions, with neutrino emission from charged pions (collisions between heavier nuclei may be approximated as bundles of nucleons that share equally the kinetic energy. See for example Breuhaus et al., 2022). Approximately 10 of the energy from the parent proton is transferred to these secondary particles, while the exact energy fraction varies with the spectral distribution of the primary protons (Kelner et al., 2006; Kafexhiu et al., 2014). The dominant gamma-ray emission mechanism in leptonic PeVatrons is IC scattering, whereby relativistic electrons up-scatter ambient photon fields, primarily diffuse infrared and the 2.7 K Cosmic Microwave Background (CMB). For electrons or positrons in the Klein-Nishina regime, the up-scattered photon energy satisfies Eγ=0.37(E/1 PeV)1.3 PeV (LHAASO Collaboration et al., 2021), implying that PeV electrons can produce gamma rays with energies 370 TeV. UHE gamma-ray emission emerges as the key observational probe for exploring PeVatron sources (see for example The HAWC Collaboration, 2023; Acero, F. et al., 2023; Angüner, 2023; Cao et al., 2023; de Oña Wilhelmi et al., 2024, for reviews.).
From the first report by the HAWC collaboration listing 3 confirmed UHE sources (Abeysekara et al., 2020), the latest catalog from the LHAASO collaboration now reports 43 UHE sources emitting above 100 TeV, all detected with statistical significance beyond 4 (Cao et al., 2024). Moreover, the firm detection of photons exceeding 1 PeV energies from the Crab Nebula (LHAASO Collaboration et al., 2021), along with the Cygnus Cocoon, a superbubble enveloping a massive SFR (Ackermann, M. et al. (2011), Fermi-LAT Collaboration; Abeysekara et al., 2021), provides proof that PeV-energy particles are accelerated somewhere in these environments (Cao & others, 2021). Regardless of whether the emission is hadronic or leptonic in origin, these findings firmly establish the Crab Nebula and Cygnus Cocoon as galactic PeVatron sources. A wide range of astrophysical objects, including SNRs (e.g. Bell et al., 2013; Celli et al., 2019; Fang et al., 2022), massive star clusters (MSCs) and superbubbles (e.g. Aharonian et al., 2019; Morlino et al., 2021; Vieu et al., 2022), PWNe (e.g. Amato et al., 2003; Ohira et al., 2018; Amato & Olmi, 2021), and microquasars (Abeysekara et al., 2018; LHAASO Collaboration, 2024c) are considered promising targets for future PeVatron searches. Additionally, all of these sources are underexplored in the UHE southern sky, where SWGO will monitor them.
The search for PeVatrons is directly connected to the long-standing mystery of the origin of CRs. The CR spectrum follows a PL distribution with an index of -2.7 above 30 GeV, transitioning at 3 PeV to a steeper slope of -3.0, forming the well-known “knee” feature (e.g. Blümer et al., 2009). Despite advances in measurements of CRs arriving at Earth up to and beyond the knee, it remains unclear which acceleration sites produce them. This limitation is well known, as particles are randomly scattering in the galactic magnetic field before arriving at Earth. In the PeV regime, it is thus necessary to look for neutral messengers such as UHE gamma-ray and neutrinos. Despite recent groundbreaking advancements in identifying PeVatrons, no galactic source has yet been unambiguously confirmed as an accelerator of hadrons to PeV energies, largely due to the difficulty in disentangling hadronic and leptonic emission scenarios. Recent studies suggest that UHE gamma-ray emission may be a universal feature of the environments of powerful pulsars, particularly those with high spin-down luminosity (e.g. Albert et al., 2021; Breuhaus et al., 2021; de Oña Wilhelmi et al., 2022), indicating that the UHE source population is dominated by PWNe. Since inverse-Compton scattering of PeV electrons on the CMB can readily produce UHE gamma rays, detecting UHE emission up to 1 PeV alone cannot conclusively establish a hadronic origin (see for example Breuhaus et al., 2021). Consequently, multiwavelength (MWL) and multi-messenger approaches become crucial for resolving this ambiguity. A clear correlation with dense target material, such as a MC, would support a hadronic scenario. However, an undeniable confirmation of hadronic PeVatron activity would come from the simultaneous co-detection of neutrinos and UHE gamma rays.
UHE gamma rays with energies beyond 200–300 TeV are also subject to attenuation through absorption, primarily due to interactions with interstellar radiation fields, particularly the CMB (e.g. Vernetto & Lipari, 2016; Zhang & Guo, 2024). This results in a reduction of UHE gamma-ray flux detected at Earth especially at these high energies, and potentially introducing artificial features into the observed spectra. Consequently, for distant sources such as the GC, these absorption effects must be corrected for an accurate interpretation of the observed spectra. The angular resolution of WCD experiments, typically around 0.2∘-0.3∘, is 3-4 times worse when compared to current IACT experiments. This reduced resolution means source confusion is an issue, making it difficult to identify the precise origin of emission through morphological analysis. Hence, combining the strengths of both methods becomes crucial when studying PeVatrons. While SWGO will be effective at identifying regions containing PeVatron activity, CTA South will offer precise localization of acceleration sites within these broader regions, allowing for detailed morphological studies.
SWGO promises significant advancements in future UHE observations (Angüner & Ergin, 2024), particularly in the southern hemisphere, which provides prime access to the Galactic Plane, as shown in Figure 1. With sensitivity comparable to LHAASO at UHE, and offering unparalleled angular resolution among wide-field instruments, SWGO has a great potential to increase the number of known UHE sources and investigate large-scale emission structures. The SWGO is expected to surpass the sensitivity of 50-hour CTA observations in detecting point-like gamma-ray sources above 10 TeV, see Figure 10. In conjunction with existing LHAASO and HAWC observations, future SWGO data will enable the first complete UHE skymap of the Galactic Plane. Consequently, SWGO is set to play a crucial complementary role in Galactic Plane surveys and to work in synergy with LHAASO, HAWC, and the future CTA project.
Beyond the GC (see Section 5.2.3 for detailed discussions), the southern sky also hosts other promising UHE sources, identified through either their high-energy fluxes or hard PL spectra without clear cutoffs, as observed by the H.E.S.S. telescopes. Among them, Westerlund 1 (Aharonian, F. et al., 2022), a young and massive stellar cluster where photons up to 81 TeV have been observed, stands out due to its extended, shell-like gamma-ray morphology spanning up to 2∘, is an ideal target for SWGO observations. Additionally, unidentified hard gamma-ray sources such as HESS J1702420A (Abdalla, H. et al., 2021), with an exceptionally hard spectral index of =1.53, and HESS J1641463 (Abramowski et al., 2014), both exhibit hard spectra and have dense target material present along the line of sight at various distances, are strong UHE source candidates potentially related to hadronic PeVatrons. Furthermore, the southern sky may also host dark UHE sources, such as LHAASO J21085157 (Cao et al., 2021b), which emits gamma rays only above 20 TeV also showing significant E100 TeV emission, and remains undetected by current-generation IACTs. Consequently, SWGO observations of the Galactic Plane from the southern hemisphere are expected to significantly expand the known UHE source population while providing crucial data above 100 TeV to enhance our understanding of the nature of galactic PeVatrons.
To investigate SWGO’s sensitivity to spectral cutoffs from point-like PeVatron sources, Monte Carlo simulations were performed using IRFs corresponding to the SWGO and SWGOA array configurations (see Figure 10 and Section 3). Gamma-ray spectra were simulated assuming an ECPL model, expressed as:
where is the normalization at reference -ray energy of E0, is the -ray spectral index, = (1/E) is the inverse cutoff energy and is the sharpness parameter describing the rate of exponential decay. This model is motivated by a parent proton spectrum that follows an ECPL form with a proton cutoff energy of E=3 PeV and =1.0, which characterizes a typical hadronic PeVatron source that contributes to observed CR spectrum, especially at energies around the knee region. The resulting gamma-ray spectrum from the pp interactions follows the ECPL shape given in Eq. 1 (see Celli et al. (2020) Eq. 17, Eq. 22 and Eq. 26). A log-likelihood ratio test, comparing a pure PL model to an ECPL model, is employed to quantify the statistical significance of gamma-ray spectral cutoffs, and defined as
where and represent the maximum likelihoods over the full parameter spaces. Spectral cutoff detection maps, following the methodology introduced and detailed in Acero, F. et al. (2023), have been produced for a broad range of ECPL flux normalization and spectral index parameters, with the cutoff energy fixed at E=200 TeV. These maps, demonstrating the sensitivity of SWGO to 200 TeV spectral cutoffs, are produced for SWGO and SWGOA array configurations, and are shown in the left and right panel of Figure 16, respectively. The figure demonstrates that the SWGO configuration exhibits strong sensitivity across a substantial range of PeVatron source parameters, even for soft sources. In contrast, the SWGOA configuration shows a considerably lower capability for detecting a spectral cutoff at 200 TeV energies, performing approximately 2530 times worse than the SWGO configuration throughout the investigated parameter space. These maps highlight the critical role of SWGO in identifying and characterizing PeVatron candidates, particularly in the southern hemisphere, where its capabilities will complement those of existing and future gamma-ray observatories. By bridging observational gaps and enhancing spectral cutoff detection at UHEs, SWGO is expected to play a key role in understanding PeV particle acceleration and significantly improve our understanding of galactic PeVatrons.
5.4 Galactic populations
5.4.1 Pulsars, PWNe, TeV halos
Pulsars are responsible for the largest fraction of identified galactic gamma-ray sources throughout the GeV-TeV range (Wakely & Horan, 2008), and are also the predominant source class associated with UHE galactic objects observed at E by LHAASO (Cao et al., 2024). Pulsars are rotating neutron stars; the pulsed radiation that gives pulsars their name originates due to the misalignment of the pulsar rotation axis and magnetic field axis.
Due to their strong magnetic field and rapid rotation, charged electrons and positrons extracted from the pulsar surface are accelerated and emit high energy photons, generating a pair-cascade process that produces copious amounts of electron-positron pairs. Under some circumstances, ions could potentially also escape the pulsar surface in non-trivial quantities (Spencer & Mitchell, 2025). Within the pulsar magnetosphere, these charged particles are tied to the magnetic field lines and co-rotate with the pulsar. At a distance from the pulsar surface at which this co-rotation would result in particles exceeding the speed of light, known as the light cylinder, the particles can escape and stream out into a magnetised pulsar wind (see Kirk et al., 2009; Philippov & Kramer, 2022, for reviews). There, they can be further accelerated at the wind termination shock (WTS), which is produced where the confining pressure of the (comparatively dense) surrounding medium reaches equilibrium with the pulsar wind’s momentum flux (Kennel & Coroniti, 1984; Nodes, 2008).
Pulsed gamma-ray emission has been detected from 294 pulsars in the GeV energy band by the Fermi-LAT in their third pulsar catalog (3PC, Smith et al. 2023), and from four pulsars at higher energies (from a few GeV up to TeV) using ground-based IACTs. Notably, the MAGIC telescopes have detected pulsed emission from the Crab pulsar as a PL tail extending from the GeV bump observed by Fermi-LAT up to energies of a few TeV (Ansoldi et al., 2016). In contrast, observations by H.E.S.S. have revealed clear pulsed emission from the Vela pulsar up to 20 TeV (Aharonian et al., 2023b), associated with the P2 pulse, characterized by a distinct spectral component separate from the GeV emission (H. E. S. S. Collaboration et al., 2023). This new spectral component is interpreted as IC scattering of electrons–responsible for the GeV emission–on soft photon fields, typically in the near-infrared to ultraviolet energy range.
To evaluate the potential of SWGO to detect pulsar TeV emission, the observed pulsed emission from Vela based on the H.E.S.S. measurements was simulated, adopting an ECPL model with cutoff arbitrarily set at TeV, given that no spectral cutoff has yet been detected by H.E.S.S. (H. E. S. S. Collaboration et al., 2023). The simulation results demonstrate that SWGO would achieve a clear detection of pulsations () within one year of data collection. Moreover, a simulated five-year exposure would allow the identification and characterization of the spectral cutoff energy. Additionally, synthetic emission featuring a softer index of 1.9, a cutoff at 20 TeV and a flux similar to the Vela pulsar was simulated. Depending on the On-phase to Off-phase ratio, such emission could be detected by SWGO in 1 to 5 years of observations.
Given these findings, SWGO will provide important observations and constraints for pulsar astronomy in the TeV regime. Its advantageous location in the southern hemisphere, grants excellent access to the Galactic Plane. In combination with SWGO’s dedicated survey capabilities, this will yield good coverage and constrain the spectral cutoff for an increasing number of TeV-emitting pulsars. This also highlights the complementarity of SWGO with other gamma-ray and MWL facilities, which in concert will provide critical insights necessary to resolve the outstanding questions outlined above.
Additionally in the magnetosphere, the magnetized pulsar wind expands into the surrounding medium resulting in the WTS, at which further particle acceleration may occur. Beyond this WTS, a nebula of energetic particles forms, predominantly electrons and positrons, that is gamma-ray bright due to IC scattering of these particles on the ambient radiation fields (Gaensler & Slane, 2006). PWNe consequently form one of the most common galactic source classes known at TeV energies, and will thus be a dominant component of the SWGO view of the Galactic Plane (H. E. S. S. Collaboration, 2018b; Cao et al., 2024).
While PWNe are believed to be leptonicly dominated, recent observations of the Crab PWN (Cao et al., 2021a) have suggested that there could be a subdominant hadronic component to the resulting gamma-ray emission at the highest energies for a small number of PWN. However, the origin of such a high-energy hadronic population is unclear, and most of the modeling work that has followed the LHAASO observations have not explained this completely (e.g. Nie et al., 2022). Models for significant hadronic acceleration in the pulsar wind have been postulated for several years (e.g. Amato et al., 2003), however this requires very high Lorentz factors for the wind in order to accelerate the particles to PeV energies. More recent work has suggested that protons may diffuse into the PWN from outside and be re-accelerated (e.g. Spencer et al., 2024). SWGO’s excellent sensitivity at high energies will help to distinguish between these possible scenarios for PWNe beyond the Crab.
Eventually, the expanding PWN may encounter a reverse shock from the progenitor SNR, causing the system to become disrupted. Such reverse shock interactions have been postulated as an efficient acceleration mechanism for CRs in older PWNe (e.g. Ohira et al., 2018). Particles escaping from the PWN following this stage may diffuse into the surrounding ISM, leading to the formation of a halo of energetic leptons, which is made visible on large scales due to further IC scattering. These pulsar halos (also known as TeV halos) were first identified in data from the HAWC observatory and have since been noted as an expected feature of PWN evolution (e.g. Giacinti et al., 2020). The distinction between PWN and halo has been a matter of debate in recent years (e.g. Giacinti et al., 2020; Sudoh et al., 2019), yet extended gamma-ray emission around energetic pulsars appears to be commonplace (e.g. H. E. S. S. Collaboration et al., 2018a; Cao et al., 2024; Albert et al., 2023, 2025).
Nearby halos, within kpc of Earth, exhibit degree-scale gamma-ray emission, enabling both the morphology and spectrum to be probed in detail. Wide-field instruments such as SWGO, and as demonstrated by HAWC, LHAASO and Tibet-AS are essential to capture the full extent of such emission. As electrons444Hereafter used to refer to both electrons and positrons unless otherwise specified. are transported away from the pulsar they undergo energy losses, leading to distinct morphological and spectral profiles that can be used to infer properties of the particle transport, such as the diffusion coefficient normalization and the level of magnetic turbulence in the region (e.g. Abeysekara, A.U. et al. [HAWC Collaboration], 2017a; López-Coto & Giacinti, 2018).
While electrons remain trapped within the PWN, the diffusion process is slow compared to average values for the ISM. However, once electrons escape during the halo-phase, it was expected that the IC gamma-ray emission from these electrons could be used to trace the diffusion within the ISM. Surprisingly, the diffusive transport remained slow even into the halo-phase (Abeysekara, A.U. et al. [HAWC Collaboration], 2017a), leading to a large number of attempts in recent years to explain this phenomenon, such as two-zone diffusion (e.g. Fang et al., 2018), CR self-induced turbulence (e.g. Evoli et al., 2018) and confinement within the region of the progenitor SNR (e.g. Fang et al., 2019). Slow diffusion also poses a challenge to local sources as an explanation for the positron excess and high energy CR electron spectrum (e.g. Aharonian et al., 2024), one that can be reconciled either by the diffusion recovering to ISM average values over time or with distance (e.g. Evoli et al., 2018) or with the additional presence of an as yet undiscovered local source (e.g. López-Coto et al., 2018). (See also López-Coto et al., 2022, and references therein.)
SWGO will contribute substantially to the study of PWNe and TeV halos by further surveying a wide variety of pulsar systems in the southern sky in an unbiased manner. Using a synthetic population of pulsars, evolved using the COMPAS population synthesis code (Riley et al., 2022), a model of PWN evolution based on H. E. S. S. Collaboration et al. (2018a) was applied to anticipate the likely gamma-ray emission along the Galactic Plane. Normalized in the northern sky to HAWC data from the 3HWC catalog (Albert et al., 2020), it is expected that SWGO can detect no fewer than PWN or TeV halo systems within 5 years of construction.
Figure 17 shows the currently known pulsar population as taken from the ATNF catalog (Manchester et al., 2005). The spin-down luminosity, , is a measure of the energy output by a pulsar over time due to loss of rotational energy and can be determined from the pulsar rotation via , where is the pulsar rotation period, the period’s time derivative and the moment of inertia. Pulsars associated with TeV emission in the vicinity are indicated by blue triangles, and those with emission reaching energies TeV indicated by orange circles. Green squares indicate pulsars for which no TeV gamma-ray emission has yet been identified. The sensitivity for a gamma-ray instrument is expected to scale as , where the relation between spin-down luminosity and anticipated gamma-ray flux is approximated using the model of H. E. S. S. Collaboration et al. (2018a). Based on this model, lines in Figure 17 indicate the anticipated sensitivity of SWGO-A to pulsar after 5 years (dashed) and for the full SWGO after 5 years (dotted). Green squares above these lines therefore indicate the discovery potential for SWGO, as well as contributions to detailed studies of other known PWNe that lie above these thresholds.
5.4.2 Supernova Remnants
SNRs play a crucial role in the evolution of galaxies and the ISM because they are responsible for enriching the surrounding space with heavy elements and accelerating CRs to very high energies. SNRs are considered the most likely source of galactic CRs, as their shocks are capable of accelerating charged particles, e.g. protons and heavier nuclei, and ejecting them into the surrounding medium. In some cases, these ejected CRs can interact with the ISM or radiation fields, leading to the production of gamma rays and neutrinos. Historically it has been considered likely that SNRs could accelerate particles up to energies around the CR knee at TeV, however the lack of SNRs seen by the LHAASO experiment at UHEs is encouraging consideration of alternative sources at the highest energies. Gamma rays produced by the interactions of these CRs with their environment enable detailed study of SNRs, providing insights into the mechanisms of CR acceleration, particle interactions, and the extreme environments within SNRs.
Theoretical predictions based on the local supernova rate estimate the number of SNRs in the Galaxy to be more than 1000. Currently, the number of detected SNRs is 300 (Ferrand & Safi-Harb, 2012). The number of SNRs detected by radio observatories has grown significantly, with over 200 SNRs identified in the southern hemisphere through dedicated radio surveys, such as the SARAO MeerKAT Galactic Plane Survey (SMGPS) (Loru, S. et al., 2024; Anderson et al., 2025, respectively).
The southern hemisphere offers a unique vantage point for observing certain SNRs, particularly those in the Galactic Plane. For example, young remnants like RX J1713.73946 (Acero et al., 2017) and Vela Jr. (H.E.S.S. Collaboration et al., 2018) are more favorably positioned for observation from the southern hemisphere, allowing for detailed observations of their gamma-ray emission. SWGO and CTAO are both planned to be in the southern hemisphere, which will offer unprecedented opportunities to study UHE gamma-ray emissions from SNRs. These observatories will be capable of resolving spatial and spectral features of these sources with higher precision and energy sensitivity than any previous instruments.
The details of acceleration at the maximum energies in SNRs is still not fully understood, and the processes that regulate escape into the ISM less so. Escaped CRs may encounter dense material such as MCs, that act as a target for interactions. The decay of neutral pions produced in these interactions can cause MCs to shine in gamma rays, making them a suitable target for SWGO providing evidence for nearby CR accelerators (e.g. Mitchell et al., 2021), and deeper insight into these poorly understood details. In many cases, MWL observations of SNRs have revealed ongoing forward shock interactions with MCs, which causes them to shine brightly in gamma rays due to the high level of CR interactions (see e.g. VERITAS Collaboration & Humensky, 2007).
To quantify the number of SNRs detectable to SWGO, the previous analysis from Scharrer et al. (2024) was repeated using the same simplistic model, but with updated sensitivity curves used throughout this work. The results of this are shown in Figure 18. With the full configuration, 9 SNRs will likely have a flux detectable within one year and be within the SWGO FoV, with another one detectable in 5 years. Roughly half of these SNRs will likely have shell-like extended emission visible to SWGO (Scharrer et al., 2024).
SWGO, in particular, will observe a larger fraction of the sky compared to previous studies of the Galactic Plane in the TeV by H.E.S.S. (H. E. S. S. Collaboration, 2018b), which will likely lead to a greater number of SNRs detected. This will constrain the acceleration mechanisms, the identification of the particle acceleration regions within the remnants providing insights into the physical conditions of these sources, and combined with multimessenger observations, offer a more complete picture of SNRs as particle accelerators.
5.4.3 Stellar Clusters and Associations
Most massive stars are formed in clusters and associations, and these objects are of long-standing interest in the acceleration of cosmic rays (e.g. Montmerle, 1979; Cesarsky & Montmerle, 1983). To date, only a few such sources have been clearly identified as TeV gamma-ray sources. However, they represent a significant scientific target for the next generation of gamma-ray observatories, especially SWGO.
In addition to the CMZ (discussed section 5.2.3), we consider two additional source types: OB associations and young massive star clusters (YMSCs). For what concerns gamma-ray emission, these can be distinguished by their compactness. Compact clusters such as Westerlund 1 contain a large number of young, massive hot stars in a compact volume, whose winds interact and drive a collective large scale wind. OB association, such as Cygnus OB2, contain a loose association of hot stars, whose winds have sufficient space to create isolated shocks and redistribute the kinetic energy before a large scale collective wind can be formed (Vieu et al., 2024). Both source types provide myriad particle acceleration sites, including the termination shocks of both individual stellar winds (e.g. Voelk & Forman, 1982) and large scale collective winds (e.g. Morlino et al., 2021), or supernova remnants. Young clusters and associations hosting massive stars will have a large number of powerful supernova events in the first several Myrs of their lives which may act as PeV CR sources (Vieu & Reville, 2023). LHAASO have reported multi-PeV photons coincident with the position of Cygnus OB2 (Lhaaso Collaboration, 2024), making such sources of great interest.
SWGO will provide an unprecedented chance to study both galactic and extragalactic (in the LMC) with improved sensitivity in the 100 TeV range. Natural candidates for observation include objects already detected in gamma rays, such as Westerlund 1 and W49A. In this regard, SWGO has a unique advantage due to its large FoV. The study of Westerlund 1 and W49A with SWGO will enhance our understanding of other well known SFRs such as the Carina nebula (H. E. S. S. Collaboration et al., 2020) or star forming regions in the LMC Aharonian et al. (2024).
1.
Westerlund 1 is one of the most massive galactic clusters. It has already been detected in the VHE range by H.E.S.S. (Abramowski, A. et al. (2012), H.E.S.S. Collaboration), but a comprehensive understanding of its spectral properties and the morphology of the region remains a challenge (see for example Härer et al., 2023). Future instruments, such as CTAO, with an improved angular resolution, are expected to be efficient in constraining the spatial origin of the emission. Additionally, SWGO is anticipated to provide valuable insights into the spectral properties, particularly in the 100 TeV range. Westerlund 1 is located close to the GC, making it a prime target for SWGO, and a challenge to observe from northern facilities.
2.
W49A, a giant MC with mass, is one of the most luminous H II regions in our galaxy. W49A hosts several active star formation sites, and is detected by H.E.S.S. with a significance of more than (Brun et al., 2010). Gas ejections and expanding shells were found in the central cluster of OB stars in W49A (Rugel et al., 2019). SWGO may help to reveal the spectral properties in the VHE and UHE gamma-ray regime.
Projections for the gamma-ray emission from stellar clusters, based on recent catalogs of data from the Gaia satellite, suggest that stellar clusters in the Milky Way could be gamma-ray bright (Mitchell et al., 2024). Such systems are good targets for wide-FoV instruments such as SWGO, due to the large anticipated size of the wind-blown bubble. Despite significant integrated gamma-ray fluxes, the large angular size of the emission will yield comparatively low surface brightness. SWGO is therefore ideal to search for such large, diffuse structures at high energies.
Following the model of Morlino et al. (2021) and Mitchell et al. (2024), Figure 19 compares predictions for the gamma-ray emission from stellar clusters listed in the Kharchenko et al. (2013) or Cantat-Gaudin et al. (2020) catalogs to the projected sensitivity for SWGO-A and the full SWGO array in 1 year and 5 years. This analytical model is subject to uncertainties of the order 10; the curves shown indicate the expected emission under the maximal scenario, a set of optimistic model assumptions, that nevertheless seem to reproduce data well in the small number of detected systems. For example, the curve shown for Westerlund 1 aligns with data from H.E.S.S. and Fermi-LAT. Contributions towards this maximal scenario include: the influence of supernovae; a high density ambient ISM; fragmentation of shell material increasing the density within the wind-blown bubble; and accounting for a systematic underestimate of the total cluster mass by a factor . We find there are at least 9 stellar clusters located in the southern sky that could be detectable by SWGO. However, Figure 19, only depicts clusters for which the angular size of the wind-blown bubble is , because as angular size increases the surface brightness of the gamma-ray emission decreases substantially, which renders the comparison to point-like sensitivity curves inaccurate. A further 3 clusters, with predicted angular sizes , have anticipated gamma-ray fluxes within the range of the clusters shown in Figure 19 and could therefore also be detectable with SWGO. One cluster in particular, NGC 6231 has a large estimated angular size (a few degrees across), yet a predicted gamma-ray flux comparable to Westerlund 1, due to its large mass and wind power comparable to NGC 3603 (Celli et al., 2024). In this case, even upper limits can place important constraints on model parameters, which provides further insights into CR acceleration in the vicinity of star clusters.
5.4.4 Binaries and Microquasars
X-ray binaries are systems in which a compact object acquires matter from a nearby companion star. Broadly speaking, they can be classified into two types: (1) Low-mass X-ray binaries (LMXBs), where the companion star has a lower mass than the compact object, and (2) High-mass X-ray binaries (HMXBs), where the compact object is the less massive component. In general, accretion in LMXBs is interpreted to occur through an accretion disk, while in HMXBs it typically results from the strong stellar winds of the companion star. However, in some HMXBs, material transfer occurs via an accretion disk, especially when the companion star has a Be spectral type, as stellar winds in these stars are not particularly strong.
Occasionally, these systems can exhibit jets of material traveling close to the speed of light. In such cases, the system is referred to as a microquasar, a name given due to its resemblance to Active Galactic Nuclei (AGN) (Mirabel & Rodríguez, 1998). A correlation has been found between the luminosity in the radio and X-ray bands and the mass of the compact object (e.g. Falcke et al., 2004), further strengthening the similarities between AGN and microquasars. Long-term radio observations of these sources have linked the presence of jets to specific accretion states. During a “hard state”—when the spectrum is dominated by energy losses through IC scattering, peaking around 100 keV—continuous outflows, known as “compact jets”, are observed in the radio band. Conversely, discrete ejections, sometimes appearing “superluminal”, are associated with the “soft state”, where the spectrum is dominated by thermal emission from the accretion disk, peaking around 1 keV. (See e.g. Remillard & McClintock, 2006; Belloni et al., 2011; Corbel et al., 2013, for further details)
If microquasars share similar mechanisms of particle acceleration and jet formation with AGN—known to be highly efficient regions for particle acceleration and to produce broadband non-thermal emission (radio to VHE gamma-ray)—it is reasonable to consider that microquasars could also act as sources of VHE gamma-ray emission. Emission in the GeV and TeV bands has been predicted for some time (e.g. Aharonian, 2004; Bosch-Ramon & Khangulyan, 2009), where it is attributed to particle acceleration within the jets. Currently, a little over 20 microquasars are known in the Galaxy (Corral-Santana et al., 2016). Only three of these have been detected at GeV energies by Fermi-LAT: Cyg X-3 (e.g. Fermi LAT Collaboration et al., 2009; Zdziarski et al., 2018), where orbital variation in gamma-ray emission has been detected; Cyg X-1, where gamma-ray emission coincides with a hard X-ray state (Zanin et al., 2016; Zdziarski et al., 2017), both classified as HMXBs; and SS 433 (Xing et al., 2019; Li et al., 2020), where variable flux was observed correlating with the precession period of the inner jet. However, the physical process driving this behavior remains unclear. In the VHE range, five microquasars have been detected: SS 433 (Abeysekara et al., 2018), V4641 Sgr (Alfaro et al., 2024c), GRS 1915+105, MAXI J1820+070 and Cygnus X-1 (LHAASO Collaboration, 2024c).
The detection of TeV emission from the jets of the microquasar SS 433 represents the first time astrophysical jets were resolved at TeV energies. TeV emission was first detected by HAWC (Abeysekara, A.U. et al. [HAWC Collaboration], 2018b), which prompted a follow-up campaign with the H.E.S.S. telescopes H. E. S. S. Collaboration (2024). The gamma-ray emission is extended and follows closely the shape of previously known X-ray structures (Geldzahler et al., 1980). Modeling of the X-ray and TeV emission suggests that it originates from efficiently accelerated electrons at the base of the extended X-ray jets (Alfaro et al., 2024b; H. E. S. S. Collaboration, 2024). More recently, the LHAASO collaboration reported emission up to 300 TeV from the SS 433 system, which might require additional hadronic contributions to the emission (LHAASO Collaboration, 2024c).
Another example is the microquasar V4641 Sgr, which was identified as a TeV gamma-ray source by HAWC (Alfaro et al., 2024c). The gamma-ray spectrum of this source is among the hardest observed, with photons detected above 200 TeV. Modeling suggests that this emission is the result of emission from high-energy protons, rather than electrons. More recently, the LHAASO collaboration reported emission up to 800 TeV from this microquasar, supporting the hadronic nature of the emission (LHAASO Collaboration, 2024c).
All the microquasar systems detected at TeV energies are located in regions of the sky visible from the Northern Hemisphere, and they were all first detected by wide-field instruments. Since the distribution of X-ray binaries does not show a preferred half of the Galaxy, it is safe to assume that this is simply because there is no unbiased gamma-ray instrument looking at the southern sky. Consequently, SWGO would be expected to at least double the number of known TeV-emitting microquasars, providing a clearer view into the population. The SWGO sensitivity is expected to peak above tens of TeV, the ideal range to probe the nature of microquasars as sources of PeV cosmic rays.
Binary stellar systems, which can be comprised of a variety of different combinations of physical components, often exhibit time-variable phenomena in their gamma-ray emission. Typically, this variation is either due to spectral modulation of the gamma-ray emission with orbital phase, or gamma-ray emission only occurring during distinct phases of the binary orbital period. An example of the former is LS 5039, a binary system comprised of a compact object and a massive O6.5V star that exhibits spectral modulation over its 3.9 day orbital period (Aharonian et al., 2006). At superior conjunction, the spectrum follows a PL shape, whereas at inferior conjunction, the flux at 110 TeV is considerably enhanced following a ECPL model. Whilst LS 5039 is observable from both the northern and southern hemispheres, several well-studied binaries are located in the southern sky.
Among these are PSR B125963 / SS 2883, a binary system with a pulsar orbiting a massive Be star (Murata et al., 2003), and eta Carinae, a colliding wind binary comprised of two massive stars, likely a luminous blue variable and a Wolf-Rayet star (Davidson & Humphreys, 1997; Iping et al., 2005). The binary pulsar PSR B125963 exhibits gamma-ray emission around the periastron passage of its 3.4 year orbit (Aharonian et al., 2005). H.E.S.S. has detected gamma-ray emission from the colliding wind binary eta Carinae around periastron in two consecutive passages of the system with 5.5 year orbital period (H. E. S. S. Collaboration et al., 2020, 2025). In this system, the gamma-ray emission is understood to arise from interactions between a stellar wind comprised of hadronic CRs and the surrounding Homunculus nebula that provides dense material for interactions during the periastron passage (Steinmassl et al., 2023).
5.4.5 Globular clusters
Globular clusters are evolved stellar systems bound by gravity, containing tens of thousands to several million stars arranged in a relatively stable and compact configuration. Due to their advanced age and the high stellar density–often exceeding 1000 stars per cubic parsec near the core (e.g. Sollima & Baumgardt, 2017)–these clusters are expected to host numerous stellar-end products, such as neutron stars and white dwarfs, many of which are trapped in compact binary systems, such as LMXBs and millisecond pulsars (MSPs; Pooley et al., 2003; Ye et al., 2019; de Menezes et al., 2023).
These characteristics make the cores of globular clusters excellent laboratories for studying the population properties of MSPs. In the gamma-ray domain, two primary spectral components are expected from these populations. The first, peaks around 3 GeV, and arises from curvature radiation in the outer magnetospheres of MSPs near their light cylinders (e.g. Caraveo, 2014; Kalapotharakos et al., 2022). The second component, peaks around 1 TeV, and originates from the continuous injection of relativistic leptons into the intracluster medium by MSPs. These leptons interact with background photon fields, such as ambient stellar photons and the CMB, producing IC emission (e.g. Venter et al., 2009; Kopp et al., 2013).
Over the past decade, the first gamma-ray component has been observed in approximately 40 globular clusters within the Milky Way using the Fermi-LAT (e.g. Abdo et al., 2009; Hui et al., 2010; Abdo et al., 2010; de Menezes et al., 2019; Abdollahi et al., 2022; de Menezes et al., 2023). These observations are generally consistent with MSP emission, although alternative explanations, including exotic scenarios such as dark matter, have also been tested (e.g. Brown et al., 2018; Evans et al., 2022). In contrast, the second gamma-ray component has yet to be conclusively detected. A notable exception may be the globular cluster Terzan 5, where the H.E.S.S. Collaboration reported gamma-ray emission coming approximately from its direction (Abramowski et al., 2011).
In this context, SWGO emerges as a powerful instrument to probe the VHE physics of globular clusters. With approximately 85% of the Milky Way’s globular clusters situated in the southern sky (Harris, 1996), SWGO’s strategic location in the southern hemisphere offers a distinct advantage over other major WCD gamma-ray instruments, which are all based in the northern hemisphere. Notably, the ten brightest globular clusters observed in the GeV range with Fermi-LAT are located in the southern sky (Abdollahi et al., 2022), positioning these clusters as prime candidates for future VHE observations and detailed study.
To evaluate the potential detectability of globular clusters with SWGO, Terzan 5 was modeled as a point-like source with a PL spectrum , with normalization cm-2 s-1 TeV-1, reference energy TeV, and spectral index , i.e. consistent with the H.E.S.S. observations from the direction of this cluster (Abramowski et al., 2011). With this spectral model, the instrument response function for the SWGO in the energy range from 0.5 TeV to 6 TeV was computed, with the pyswgo software package555https://gitlab.com/swgo-collaboration/irf-production. The results demonstrate that over a 1 year observation period, SWGO could achieve a detection significance of nearly . This is a significant improvement over the achieved by H.E.S.S. with 90 hours of observations using 3- and 4-telescope combinations. The superior performance of SWGO is primarily attributed to its enhanced sensitivity to photons with energies above 1 TeV.
The main results expected to be achieved with SWGO on the topic of globular clusters are listed below.
1.
Extended gamma-ray emission: in the GeV energy range, globular clusters are expected to appear as point-like sources due to the concentration of MSPs in the cluster cores, a consequence of dynamical friction (de Menezes et al., 2023). However, in the VHE domain, the situation changes. Relativistic leptons that escape the magnetospheres of MSPs can propagate within the cluster, and scatter soft stellar photons to energies that exceed 100 GeV through the IC process (e.g. Bednarek & Sitarek, 2007; Bednarek et al., 2016). Depending on the propagation behavior of these leptons, the associated IC emission could extend beyond pc from the cluster centers (e.g. Cheng et al., 2010), and potentially produce extended gamma-ray sources detectable within SWGO’s angular resolution.
2.
The correlation between gamma-ray luminosity () and stellar encounter rate () is well established in the GeV range, where the gamma-ray luminosity of globular clusters scales linearly with their stellar encounter rate (Abdo et al., 2010; Hui et al., 2010; de Menezes et al., 2019). Using this correlation, one can estimate the total number of MSPs in a cluster independently of radio observations. One major issue is that in the GeV band, the gamma-ray emission of MSPs is beamed, leading to significant scatter in the correlation. de Menezes et al. (2023) show that this problem is mitigated in X-rays, where the emission from compact objects is expected to be roughly isotropic, and this should also be the case for the VHE range, where gamma-rays are roughly isotropically scattered following lepton propagation within the cluster. Consequently, we anticipate a much tighter correlation in the VHE domain compared to the GeV range, enabling more precise estimates of the total MSP population in globular clusters. Furthermore, de Menezes et al. (2023) shows that this linear correlation holds only for clusters with high stellar encounter rates (, normalized with respect to 47 Tucanae, which has ), which suggests that channels other than stellar encounters can form MSPs in globular clusters. If a similar correlation were observed in the VHE regime, it would provide valuable insights into this phenomenon, and help to constrain the underlying mechanisms driving the relationship between and .
3.
Magnetic field, lepton injection spectra, and diffusion coefficient: the internal physical conditions within globular clusters remain poorly understood. A direct measurement of their VHE gamma-ray spectra could provide critical insights into key parameters, including the average magnetic field strength, the energy spectra of injected relativistic leptons, and the mechanisms governing CR propagation in these environments (e.g. Kopp et al., 2013; Bednarek et al., 2016; Ndiyavala et al., 2018). Such measurements would represent a significant step toward unraveling the complex interplay of astrophysical processes within globular clusters.
4.
Non-accreting white dwarfs: the escape of relativistic leptons from the magnetospheres of non-accreting white dwarfs in globular clusters may also upscatter surrounding soft photons to TeV energies via the IC process (e.g. Bednarek, 2012). While the spin-down luminosities of white dwarfs are typically about 1000 times lower than those of MSPs, they may outnumber MSPs by a similar factor. Consequently, white dwarfs could contribute significantly to the observed TeV emission from globular clusters. Observations with SWGO may provide the means to disentangle the contributions of white dwarfs and MSPs, potentially via the identification of distinct spectral components in the VHE gamma-ray spectra of these clusters. Such a breakthrough would enable detailed studies of both source populations, and offer new perspectives on their roles in the astrophysical processes shaping globular clusters.
Beyond Terzan 5, several other globular clusters are promising targets for SWGO, with predicted VHE gamma-ray emissions comparable to, or potentially exceeding, that of Terzan 5. Notable examples include NGC 6388 and NGC 362, identified in previous studies (see Ndiyavala et al., 2018, for a comprehensive list).
5.4.6 Novae
Nova explosions occur essentially at random across the galaxy, so SWGO’s the wide FoV will increase the likelihood of rapidly obtaining exposure on novae. A nova happens when a white dwarf accreting matter from its companion star (i.e. a white dwarf-star system), accumulates enough hydrogen on its surface to produce a thermonuclear explosion (e.g. Chomiuk et al., 2021). One of the best examples for TeV gamma-ray emission from such a system was RS Ophiuchi (RS Oph), which was detected by optical telescopes, Fermi-LAT in GeV energies, as well as H.E.S.S. and MAGIC in TeV energies, simultaneously (Acciari, V.A. et al. (2022), MAGIC Collaboration; Aharonian, F. et al. (2022), H.E.S.S. Collaboration). Since a large portion of the Galactic Plane will be visible to SWGO, it is anticipated that a number of nova explosions may occur within the FoV of SWGO, which will enable prompt observations and/or constrain limits. This will be useful to understand the underlying physical conditions leading to TeV gamma-ray emission from white dwarf-star systems (e.g. Zheng et al., 2024).
Nova events are typically first identified due to a strong increase in optical brightness over a few hours. Although their optical light curves can either decay smoothly or exhibit variation such as jitters, dust dips or oscillatory behavior, they typically return to optical quiescence after year (e..g. Strope et al., 2010). By contrast, the gamma-ray emission detected by Fermi-LAT generally ceases to be detectable within days (e.g. Ackermann et al., 2014b). Indeed, the only TeV-detected nova to date, RS Ophiuchi, remained detectable for weeks, however the emission was much reduced after the first few days (Aharonian, F. et al. (2022), H.E.S.S. Collaboration).
The sensitivity of the full SWGO array over a period of 30 days was evaluated, and an integral sensitivity of in the energy range 0.5 to 100 TeV was obtained. To compare this to the anticipated gamma-ray energy flux due to hadronic emission from a nova, the model of Acciari, V.A. et al. (2022) (MAGIC Collaboration) was adopted, which enables an evaluation of the gamma-ray energy flux as a function of distance and shock velocity. Thus, for the closest likely novae such as T CrB, at a distance of pc, a shock velocity of order km/s is required for detection by SWGO. At a distance more comparable to RS Oph (2.5 kpc), a much higher velocity km/s is required, which is somewhat higher than the km/s that was observed during the 2021 outburst. Therefore only a small number of truly exceptional nova events are likely to be detectable with SWGO. Nevertheless, due to its large FoV and high duty cycle, SWGO remains poised to provide data coincident with novae more systematically than has been feasible with IACTs. For example, while the full Moon caused the cessation of IACT observations during the 2021 RS Oph outburst, such an occurance would have no effect on SWGO data acquisition (Aharonian, F. et al. (2022), H.E.S.S. Collaboration; Acciari, V.A. et al. (2022), MAGIC Collaboration).
6 Transients and Variable Sources
SWGO will provide extensive sky coverage, which will enable the continuous monitoring and real-time detection of highly variable emission. This includes not only AGN flares, but also transient phenomena such as Gamma-Ray Bursts (GRBs), which can occur anywhere in the sky without prior warning. With its wide FoV, SWGO will be capable of recording sudden increases in gamma-ray flux across multiple regions of the sky simultaneously, without the need to re-point at a specific source. This capability will allow for a rapid response to transient events, and offers a significant advantage for continuous sky monitoring. Alerts generated by SWGO are likely to be critical for pointed instruments to obtain information from the early phase of transient events.
In addition to AGN and GRBs, there is potential to deeply probe transients which are so far not established as VHE-UHE gamma-ray source, for example Fast Radio Bursts (FRBs) or luminous fast blue optical transients (LFBOTs). In all cases, measurements in the gamma-ray band will complement those across the electromagnetic spectrum in constraining the physical processes at work.
Extragalactic sources are attenuated due to the extragalactic background light (EBL), which limits the energy range, and/or distance range, at which they can be observed. The expected sensitivity of SWGO at very low energies, from 100 to 300 GeV, will be an order of magnitude better than current wide FoV detectors, which makes it a highly relevant instrument to locate and monitor extragalactic transient sources, as well as to provide rapid alerts to trigger complementary studies with other present and future instruments, such as the CTAO.
6.1 Active Galactic Nuclei
Active galaxies contain a nuclear region at least 100 times brighter than their integrated starlight due to the accretion of matter onto a supermassive black hole (SMBH) with masses up to (e.g. Blandford & Rees, 1978; Urry & Padovani, 1995). There are several classification schemes for active galaxies, but only jetted AGN have been observed up to TeV energies. Regardless of orientation, most of these AGN show rapid flux variability at very high energies (VHE), suggesting that the emission occurs from small regions. The exact location of the VHE gamma-ray emission is uncertain, but it is thought to originate either along the jet (e.g. Rieger & Levinson, 2018) or at its base, near the SMBH (e.g. Raue et al., 2012). The nearby active galaxy Cen A is a notable exception - with TeV emission resolved along the kpc-scale jets H. E. S. S. Collaboration (2020).
The mechanisms responsible for particle acceleration in AGN remain a significant open question. Fermi shock acceleration is often assumed to contribute, as it is a well established mechanism to produce non-thermal particle spectra over many decades in energy. Observational support is provided by the detection of shocks in the radio spectra, which occur further out along the jet. Second-order Fermi acceleration due to MHD turbulence, as well as shear acceleration have been proposed to explain the smooth evolution of the spectral index (e.g. Rieger et al., 2007; Wang et al., 2021). Magnetic reconnection has also been suggested as a possible acceleration mechanism that may account for observed phenomena, especially rapid variability and large energy outputs (e.g. Giannios, 2010; Sironi et al., 2015; Petropoulou et al., 2016).
The difference between blazars and radio galaxies is the angle between their jet axis and our line of sight. Blazars are observed along the jet axis, and dominate the VHE sky outside the GP. They account for 70 of the extragalactic sources detected at VHE energies (Cerruti, 2020). Radio galaxies are observed at larger angles, which allow for the study of the most inner region of the AGN. At the moment there are four radio galaxies detected at VHE gamma-rays, including M 87, Cen A, NGC 1275 and 3C 264 (Rulten, 2022), with only Cen A located in the Southern sky, and M 87 visible from some Southern Hemisphere locations.
The broadband emission from blazars and radio galaxies is characterized by two dominant components in their spectral energy distribution (SED), which are often attributed to electron synchrotron at lower frequencies and IC at higher frequencies, although the identity of the high-frequency peak is a matter of some debate. Blazars can be further classified into flat spectrum radio quasars, Low-synchrotron-peaked BL Lacs (LBL), Intermediate-synchrotron-peaked BL Lacs (IBL), High-synchrotron-peaked BL Lacs (HBL) and Extreme high-synchrotron-peaked BL Lacs (EHBL) depending on the peak frequency of the electron synchrotron component of the SED. Some HBL sources can become EHBL when in a flare state, and these have been observed to reach up to tens of TeV.
6.1.1 Emission Scenarios
AGN can exhibit MWL emission from radio up to VHE gamma-rays. As stated previously, the SED of AGNs is composed of two primary components. The low-energy component is well understood as the result of synchrotron emission from relativistic electrons. The second component, responsible for gamma-ray emission from MeV up to TeV of energy, is usually but not universally attributed to inverse Compton emission from the same electrons, either in external radiation fields (e.g. Dermer et al., 1992; Sikora et al., 1994), or via the synchrotron self-Compton (SSC) mechanism (e.g. Maraschi et al., 1992; Bloom & Marscher, 1996). Whilst single zone models have been very commonly applied in the past, it is clear that these are complex objects with multiple-zone models increasing invoked (e.g. Kapanadze et al., 2018; Sahu et al., 2021). Hadronic emission is sometimes invoked either for the entire high-energy peak or in combination with IC emission (e.g. Mannheim & Biermann, 1989; Mastichiadis & Kirk, 1995; Mücke & Protheroe, 2001; Fraija, 2014). Continuous monitoring of AGN during both quiescent and flare states, across the electromagnetic spectrum, is crucial for constraining these models. Note that whilst VHE gamma-ray activity is often correlated with that seen in the optical and X-ray, blazars have also been reported to exhibit orphan flares in the TeV band (Krawczynski et al., 2004).
The VHE horizon is limited by gamma-ray interactions with photons of the diffuse (FIR-UV) EBL, such that most AGN detected by SWGO are expected to be located at moderate redshift. Exceptions to this are powerful flaring episodes of more distant objects, anomalous optical depths due e.g. to Axion-Like Particles (see Section 7.3), or production of gamma-rays from secondary particles closer to the observer (e.g. Takami et al., 2013).
6.1.2 AGN in the SWGO FoV
Figure 20 shows the AGN that could be observed by SWGO-A and SWGO after 1 and 5 years of observation. The sources shown are taken from the Fermi-LAT 4th AGN catalog (Ajello et al., 2022), with their fluxes extrapolated to VHEs and limited to redshifts z0.4. The left panel presents the extrapolated fluxes and accounts for attenuation due to the EBL, while the right panel includes both EBL attenuation and an additional exponential cutoff at 1 TeV. This cutoff is introduced to avoid an overestimate of the energy flux of each source, and thus the number of sources potentially detectable. Under the more optimistic scenario in the left panel, SWGO-A would be able to observe 4 sources after 5 years, and SWGO would detect approximately 53 sources after 1 year. In the more conservative scenario, shown in the right panel, SWGO would detect about 15 sources over a 5 year period. These estimates do not include the potential for flare activity, which may increase the total number of detectable AGN, but not provide consistent coverage of them.
SWGO will also be able to observe different activity states of AGNs and potentially monitor their flare episodes. For example, in the case of the HBL blazar PKS 2155304, located at z=0.116, both SWGO-A and SWGO will be able to detect the source during its high and low states (see Figure 21).
Whilst the number of detected objects is modest in comparison to CTAO, the SWGO measurements will be important in terms of understanding the long-term average behavior of sources, probing highest energy emission and providing triggers to pointed instruments like CTAO.
6.2 Gamma-Ray Bursts
GRBs are among the brightest events in the universe. Their emission is split into two phases: a highly variable prompt phase (from optical to MeV energies) and a longer-lasting afterglow across all wavelengths. Prompt emission durations follow a bimodal distribution, dividing GRBs into short (2 s) and long (2 s) types. Long GRBs are commonly associated with massive star collapses, while short GRBs are typically linked to compact object mergers (e.g. Duncan & Thompson, 1992), although exceptions and alternative scenarios remain under investigation.
The emission in both the prompt and afterglow phases are thought to arise from synchrotron and SSC processes: synchrotron can reach up to tens of GeV, while SSC dominates from tens of GeV to TeV energies.
Until recently, GRB afterglows were only observed up to approximately one hundred GeV. However, detections from GRBs 180720B, 190114C, 190829A, and the exceptionally bright GRB 221009A (the Brightest Of All Time, BOAT) have revealed photons up to nearly 4 TeV—and beyond 10 TeV in the last case, with VHE emission lasting thousands of seconds post-trigger (e.g. Abdalla et al., 2019; MAGIC Collaboration et al., 2019; H.E.S.S. Collaboration et al., 2021; Cao et al., 2023a). These discoveries confirm that GRB afterglows can emit VHE gamma rays for several days, challenging current models.
So far, VHE emission has only been firmly detected from long GRBs, but a detection in the VHE range would place powerful constraints on the physic models of short GRBs from compact binary mergers. SWGO-A, with its broad sky coverage and rapid localization (<10 arcmin within seconds), can greatly enhance MWL follow-up, especially crucial for GW-related events.
While prompt VHE emission remains undetected, and remarkably absent in the BOAT, SWGO-A is well-positioned to investigate this open question. The next sections, first focus on a detailed case study of GRB 190114C, with simulations of early-time emission, with and without prompt VHE flux, to illustrate how SWGO-A could complement CTAO-South in specific, well-characterized events. We then explore SWGO-A’s detection prospects using a synthetic population of GRBs, which provides a statistical view of its sensitivity across various scenarios.
6.2.1 GRB 190114C as a Case Study
Taking the measured light curve of GRB190114C (MAGIC Collaboration et al., 2019) two scenarios are considered for the early development of the GRB light curve at VHEs: with and without prompt phase emission. This flux evolution is simulated as would be observed with SWGO-A at a zenith angle of 20∘. The spectral and temporal properties are selected to match those of GRB190114C, with a spectral index of 2.2 and temporal evolution . To explore a domain more relevant to GW-detected mergers, a smaller redshift of is adopted for this simulation. Note that in reducing the GRB’s redshift, the event becomes less luminous, as the (unabsorbed) flux normalization was kept to be the same. The effect of EBL absorption is accounted for with the model proposed in Saldana-Lopez et al. (2021). The prompt phase lasts 25 s (analogus with GRB190114C) and is simulated as either no emission followed by a rapid rise or a flat plateau. To illustrate the complementarity of the (almost) co-located CTAO-South the GRB is considered to be observed with the CTAO-South configuration with a delay of 60 s. Observations with four Large-Sized Telescopes (LSTs; the largest component telescopes of CTAO) are also considered, which have a more rapid follow-up response, reducing the delay to 40 s. The results are shown in Figure 22 and highlight how even the rapid response of the LSTs misses the prompt phase, such that only SWGO (and SWGO-A) can distinguish between the two scenarios.
6.2.2 SWGO GRB Detection Rates
To perform the detectability study for SWGO, presented here, a synthetic population of GRBs was simulated with the Conditional Tabular Generative Adversarial Networks (CTGANs, Xu et al., 2019). A CTGAN is a specific type of GAN, or machine learning method in which a generator and a discriminator network compete to create increasingly realistic data. CTGANS are designed explicitly for synthesizing realistic tabular data by learning conditional relationships between features. To apply CTGANS to GRBs, the GRB data from Zhu et al. (2023) and Tang et al. (2019) that contains a clean sample of short and long GRBs was used, respectively, with parameters such as redshift, duration (T90), fluence, and isotropic energy (Eiso) in a tabulated form. This GRB list was cross-matched with the Swift GRB table (NASA Goddard Space Flight Center, 2025) to include the BAT photon index and XRT 11 hrs flux. The feature engineering process before training the CTGAN applies a logarithmic transformation to these variables. For each column, the standardized (zero-mean, unit-variance) version of that column (z-score) was added. To improve generalization and reduce overfitting, a multiplicative Gaussian noise function was used for both the logarithmic and linear values. An isolation forest with contamination set to 0.05 was used, to ensure that extreme outliers do not bias the generative model. The model was trained with the ctgan.CTGAN library of the Synthetic Data Vault Project, and the tuned parameters include: epochs, batch size, generator and discriminator learning rate, and pac.
After cleaning, 80% of the data were used to train the model with the remaining 20% held for validation. The evaluation metric employed to select the best model was the LogisticDetection, which picks the model with the lowest value. With this best model, a sample of thousands of GRBs was created to compare features like Eiso, T90, fluence and the relations between the input/real data and the synthetic one. This pipeline ensures the synthetic GRB population is physically meaningful and statistically consistent. The final GRB population sampled 10 000 GRBs with a ratio of 1/4 between short and long GRBs. This number of GRBs is equivalent to approximately 100 years of Swift observations, since it detects roughly 100 GRBs per year (NASA Goddard Space Flight Center, 2024). The distribution of GRBs in the sky was randomly sampled within the Swift range of (-20∘ , 70∘ ) in declination, with random values for the right ascension coordinates and onset times. Although Swift has detected GRBs across the whole sky, the declination region considered encompasses the majority of its cumulative sky exposure and GRB detections (90%), while excluding zones with limited visibility due to Sun/Earth avoidance constraints. This conservative cut ensures uniformity of detection efficiency without a requirement to model detailed exposure maps.
The synthetic GRB population was injected into the SWGO simulation pipeline, randomly distributing the bursts in time. A power-law spectral model was assumed, with an index of 2, and normalized under the assumption that the fluence per decade remains constant. The effect of EBL absorption was taken into account for the redshift of each GRB. For the temporal evolution, a linear rise up to T90 was assumed, and followed by a power-law . For each event, the detection significance was derived by integrating within a window of length 3 T90. Note that this integration window is chosen for simplicity, and could be further optimized depending on the brightness of the GRB.
The random distribution of bursts in time translates to a random distribution of sky coordinates. The position of the GRB is tracked within the integration window in steps of 1 min to account for the changing response as a function of zenith angle. The instrument response is simulated for the zenith range 0∘ -52∘ , so emission from the GRBs is only simulated while they fall within this range. The simulation was run 200 times, each with a different random seed for the simulated counts and burst time assignment. An example realization of the simulation and the cumulative yearly rate of detected GRBs are shown in Figure 23. From the right panel in the figure it can be seen that the total rate of GRBs detected by SWGO is estimated as 0.7 events per year from externally triggered GRBs or those with X-ray counterparts. Looking ahead, the number of GRB triggers is expected to increase with the operation of recent missions such as the Space Variable Objects Monitor (SVOM), launched in mid-2024, and the Einstein Probe, launched in early 2024, both of which have been successfully commissioned. The inclusion of these additional triggers could potentially increase the SWGO detection rate to 1.4 GRBs per year.
The GRB detection prospects study for SWGO assumes that only GRBs which can be associated to an X-ray/soft gamma-ray trigger. Blind searches are of considerable interest, but will contribute at a lower rate to the detection of SWGO GRBs.
6.3 Exploratory Searches for New Transient Phenomena
The Universe abounds with transient phenomena that reveal extreme and energetic processes. From the explosive deaths of massive stars to the violent interactions in binary systems, these fleeting events provide a unique opportunity to understand the physics of compact objects, particle acceleration, and the behavior of matter under extreme conditions.
Galactic and extragalactic transients, in particular, offer a rich laboratory for studying the diverse and complex processes at play in these extreme environments. While phenomena such as GRBs, AGN, X-ray binaries, and microquasars have been extensively studied, there remains a vast landscape of poorly explored transient events. Among them there are a large set of galactic transients, such as magnetar outbursts, flares from young stellar objects, and tidal disruption events near the Galactic Center, among others.
Magnetar bursts are extremely energetic events associated with highly magnetized neutron stars, which possess magnetic fields on the order of to Gauss (e.g. Kaspi & Beloborodov, 2017). These objects can be classified as Soft Gamma Repeaters (SGRs) or Anomalous X-ray Pulsars (AXPs), and produce bursts of gamma rays and X-rays due to magnetic reconnection in their crust or the release of energy stored in their magnetic fields (e.g. Thompson & Duncan, 1995). Theoretically, it is believed that instabilities in the neutron star’s crust, combined with the twisting and breaking of magnetic field lines, accelerating particles that can generate high-energy emission, including gamma rays (e.g. Harding & Lai, 2006). To date, approximately 30 magnetars have been identified in our galaxy, both as SGRs and AXPs, with some of the most studied being SGR 1806-20 and SGR 1900+14, known for their giant bursts and recurrent flares (Esposito et al., 2021; Palmer et al., 2005).
Analyses of the current transient source catalog reveal SWGO’s significant potential for monitoring both galactic and extragalactic phenomena. This study shows that among the 166 confirmed galactic sources, 140 fall within SWGO’s optimal viewing window, spanning various object classes including binaries, black holes, and neutron stars. The instrument’s capability to monitor high-energy sources is particularly noteworthy, with coverage of 6 high-energy binary systems, 4 black holes, and 3 pulsars. Of special interest are the rare galactic events showing multi-messenger signatures, including novae and SNRs, which represent prime targets for SWGO’s transient monitoring program. The catalog also includes 6 SGRs, with one confirmed high-energy emitter, underlining SWGO’s potential for magnetar studies. A broader analysis of all transient alerts reveals that 22.2% of reported events fall within SWGO’s visible region, with an additional 4.7% in challenging but potentially observable zones. This comprehensive coverage, combined with SWGO’s continuous monitoring capability, positions the observatory as a crucial facility for both targeted observations and serendipitous discoveries of transient phenomena. While classification efforts are ongoing, with about 54% of historical alerts still awaiting coordinate confirmation, the current statistics demonstrate SWGO’s capability to significantly contribute to our understanding of the dynamic gamma-ray sky.
The large sky coverage and high duty cycle of SWGO is ideal for an all-sky search of transient and variable phenomena in the extragalactic TeV gamma-ray sky. With its low energy threshold, SWGO can conduct complimentary observations to other experiments such as Fermi-LAT and search for variability of AGN, as already seen in the Fermi-LAT All-sky Variability catalogs, 1FAV and 2FAV (Ackermann et al., 2013; Abdollahi et al., 2017, respectively).
7 Particle Physics and Beyond the Standard Model
This section explores how SWGO’s unique combination of wide-field, wide-energy range performance and southern hemisphere location will contribute to our knowledge of particle and fundamental physics.
This includes probing Lorenz Invariance, testing various dark matter (DM) particle candidates including WIMPs and primordial black holes (PBHs), and in general probing subtle signatures predicted by beyond-the-standard-model (BSM) theories.
7.1 Dark Matter (DM)
Among various BSM scenarios, the nature of DM remains one of the most compelling and accessible to experimental investigation. Although evidence for DM spans all scales of the cosmos, its elusive nature and apparent lack of detectable signals remain a major unsolved challenge in modern astrophysics. One of the most extensively studied classes of DM candidates is that of weakly interacting massive particles (WIMPs). These are hypothesized to have masses in the GeV-TeV range and interact through gravitational and weak forces. In this mass range, the annihilation or decay of WIMPs into secondary particles, such as gamma rays, is predicted to produce potentially detectable gamma-ray signals in astrophysical sources. These signals are theorized to peak in regions with high DM density, such as the GC and nearby satellite galaxies. Given the coverage of Southern sky (with prime access to these regions), continuous monitoring and wide FoV and high sensitivity (giving it an edge over Southern IACTs like H.E.S.S. and CTAO-South), DM detection is one of the primary scientific goals of SWGO.
SWGO’s DM searches will be highly complementary to those of other experiments, such as Fermi-LAT and CTAO. The combined data from these experiments could place stringent constraints on the thermal relic annihilation cross-section of WIMPs with masses up to 100 TeV (see Figure 24, right panel, and McDaniel et al. (2024)). Particularly, in the mass range from a few hundred GeV to few tens of TeV, joint observations of the GC by CTAO and SWGO could result in coincident detections, thereby significantly improving its statistical significance. In this complementary approach, SWGO measurements are expected to probe the spectral cut-off while CTAO observations would be essential to resolving the spatial morphology of the signal (see Figure 24, left panel).
7.1.1 Sensitivity to DM Annihilation and Decay
To quantify the observational prospects for SWGO, we begin by expressing the expected gamma-ray flux from DM self-annihilation, , and decays as
and
where on the right side of equality, the first parenthesis denotes the particle physics term and second the astrophysics term.
The particle physics terms contain the DM particle mass, , the velocity-weighted annihilation cross-section, , DM lifetime, , and the differential spectrum of gamma rays in a specific annihilation or decay channel, .
The astrophysical factors, also called J-factor for annihilation reactions and D-factor for decays, are defined as
and
where is the DM density distribution. Both astrophysical factors are expressed as integrals along the line of sight (l.o.s.) and the solid angle, over for annihilation and for decay. In this work, we restrict our analysis to the case of self-annihilating dark matter; the scenario of decaying dark matter will be explored in a future publication.
For up to 100 TeV, we use the spectra provided by the PPPC4DMID (Cirelli et al., 2011), which is widely adopted in the community for indirect detection analyses. Since PPPC4DMID does not include predictions for masses above 100 TeV, we employ the HDMSpectra package (Bauer et al., 2021) in that range, thereby extending the computation of DM-induced gamma-ray spectra into the multi-PeV range and up to the Planck scale. However, it is important to note that in the higher mass range, most DM models will be in violation of unitarity bounds (Smirnov & Beacom, 2019; Griest & Kamionkowski, 1990) (for some exceptions, see e.g. Berlin et al. (2016a, b)).
To analyze the expected signal against the background, we use a 2D (spatial and energy) binned joint-likelihood method. Assuming a Poissonian distribution for the detected events, likelihood functions are calculated for each bin and combined to form a joint likelihood for the entire dataset. Using the combined likelihood, we can search for a signal or derive limits to the parameter space of the dark matter particle at different confidence levels (Viana et al., 2019).
7.1.2 DM Annihilation Searches towards the Galactic Halo
Given its high DM density, the GC is expected to yield the strongest potential signal from DM annihilation or decay, making it a prime target for either detecting a signal or setting upper limits on the annihilation cross-section or decay lifetime. However, the signal is heavily contaminated by astrophysical background, making the derived limits highly dependent on the assumptions about these astrophysical sources. To circumvent this issue, we exclude the galactic plane from our analysis and focus primarily on searching for signals within the Galactic halo. In the calculations presented here for the GC, the spatial binning uses a bin size of 0.3° for a 6°6° square region centered on the GC, with a mask on the galactic plane of 0.3°. Energy binning consists of 25 logarithmically-spaced bins between 31.6 GeV and 2 PeV. In order to compare SWGO sensitivity limits to other observatories, we assume a common Einasto profile (Acharyya et al., 2021; Abdalla et al., 2022) for the DM density distribution of the galaxy.
Figure 24, right panel, shows the 95% C.L. sensitivity on versus for 1 and 5 years of observation by SWGO-A and SWGO in the case of DM particles annihilating and producing a pair. This choice of annihilation channel is conservative, as other channels (e.g., or ) tend to yield harder gamma-ray spectra and thus stronger signals at the high-energy end of the spectrum (Viana et al., 2019). The sensitivity of CTAO (for the case of only misidentified CRs and with a galactic plane mask on ) (Acharyya et al., 2021), HAWC (Albert et al., 2023b), and H.E.S.S. (Abdalla et al., 2022) are plotted for comparison. A sensitivity smaller than the thermal relic cross-section cm3 s-1 is reachable for SWGO in the mass range of 800 GeV to 140 TeV. The sensitivity is better than that of CTAO for all masses 5 TeV and several orders of magnitude better than that of HAWC. Within 5 years of operation, SWGO-A will already have sensitivities comparable to those of H.E.S.S., obtained after two decades of operation and currently the most sensitive instrument in the TeV mass domain.
7.1.3 DM Searches in Satellite Galaxies
Dwarf spheroidal galaxies (dSphs) are among the most DM-dominated objects known and are sufficiently close to Earth to potentially produce detectable signals. Their appeal as targets for DM searches stems from two key factors: their exceptionally low levels of astrophysical background, often referred to as being “background-free”, and their relative proximity to Earth. The low background is attributed to their high mass-to-light ratios and lack of significant star formation or other astrophysical processes. These characteristics make dSphs prime candidates for observing gamma-ray signals from DM annihilation or decay.
In the calculations presented here, similar to those in Section 7.1.2, the energy binning consists of 25 logarithmically-spaced bins between 31.6 GeV and 2 PeV. In addition, the joint-likelihood method is used to combine data from 32 distinct dSphs, rather than analyzing different spatial bins. These dSphs were selected from a list of 50 that were recently studied by Fermi-LAT (McDaniel et al., 2024), as they lie within the FoV of SWGO. They include dSphs with high measured J-factors, such as Coma Berenices ( ), Reticulum-II () and Horologium-I (), as well as dSphs with high estimated J-factors, such as Cetus-II () and Carina-III (). The galaxies which dominate the combined sensitivity are determined by a combination between high J-factor and large yearly exposure to SWGO.
Figure 25 (left panel) shows the 95 % confidence level (C.L.) sensitivity on as a function of for 1 and 5 years of observation with SWGO-A and SWGO, assuming DM particles annihilate into . We compare our results with the Coma Berenices limit from CTAO (Saturni et al., 2023), the combined limit from HAWC (Albert et al., 2020), the combined limit from Fermi-LAT (McDaniel et al., 2024), and the combined limit from LHAASO (Cao et al., 2024). Notably, SWGO exhibits significant improvements in sensitivity, surpassing the capabilities of current gamma-ray instruments such as HAWC, primarily due to its superior flux sensitivity. Moreover, in the multi-TeV DM mass range, SWGO is projected to achieve better sensitivity than CTAO due to its higher exposure and simultaneous coverage of multiple targets.
Due to its continuous monitoring capabilities, SWGO can accumulate data across all regions of the Southern gamma-ray sky, including those where new dSphs may later be discovered, unlike IACTs, which require targeted observations. For instance, the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019) is expected to discover hundreds of new dSphs in the Southern Hemisphere with unprecedented sensitivity (Manwadkar & Kravtsov, 2022). Once identified, these new targets can be promptly analyzed using existing SWGO data, offering a major advantage for indirect DM searches.
The right panel of Figure 25 illustrates the projected improvement in the sensitivities of SWGO and SWGO-A to DM annihilation cross-sections based on the anticipated increase in the number of known dSphs. Assuming the J-factor distributions of newly discovered dSphs are similar to those currently known and that they are isotropically distributed across the sky, an improvement in sensitivity by a factor is expected.
7.2 Primordial Black Holes (PBHs)
Beyond particle DM candidates, SWGO may also probe early-universe relics such as PBHs, which could contribute to the DM abundance. Black holes (BHs) are solutions of Einstein’s field equations for a point-like source of mass . The simplest solution is for the Schwarzschild metric (spherical symmetry), which defines the event horizon at a radius, , where and denote the gravitational constant and the speed of light, respectively. BHs resulting from stellar gravitational collapse span a mass range from a few to millions of solar masses. However, in early stages of the universe during the radiation-dominated era, the collapse of over-densities resulting from strong inhomogeneities may have led to the formation of low-mass BHs, known as PBHs. Furthermore, this hypothetical mechanism may extend the BH mass spectrum up to the Planck scale, g (Hawking, 1971). Since PBHs are non-baryonic, they are considered viable candidates for cold DM, attracting increasing interest as partial contributors to (or even the main constituents of) DM in the universe (Carr et al., 2016).
In the 1970s, S. Hawking applied quantum field theory to spacetime in the vicinity of the event horizon, concluding that BHs can radiate particles of any spin that follow a thermal spectrum with temperature g GeV, where is the reduced Planck constant and the Boltzmann constant (Hawking, 1974). As the BH increases in temperature, heavier particles start to be produced, that decrease the BH mass at a rate , where accounts for the degrees of freedom of the emitted particles. Integrating the mass-loss rate gives the initial mass, , required to be reduced to zero (evaporation) in time, , according to the relation g, where the value of is from Ukwatta et al. (2016), which considers only Standard Model (SM) particles. PBHs with initial masses g, should have comparable to the age of the Universe and, therefore, be evaporating today.
PBH evaporation ends with an energetic short-lived particle burst. The energy of the emitted gamma rays depends on the physics that govern the last stages, i.e., the number of degrees of freedom of the particles that are being emitted. In the Standard Model evaporation scenario, the energy of gamma rays in the final explosion covers the range from MeV to TeV. In particular, photons in the 1-100 TeV band are emitted for evaporation times to a few seconds (Halzen et al., 1991). Gamma-ray emission primarily occurs through two channels: (a) photons from Hawking radiation and (b) photons from hadronic decays during quark and gluon fragmentation (when the temperature of the PBH reaches - MeV). In the latter, the neutral-pion decay into two photons is the major contribution to the spectrum. For a PBH with evaporation time and temperature at the beginning of the evaporation time , the number of photons per unit energy can be parametrized as (Petkov et al., 2008)
for photons with GeV. Although the final stages of PBH evaporation may resemble short GRBs, their light-curve exhibits some distinctive features (Boluna et al., 2024):
1.
Short duration: The burst typically lasts from a few seconds to several tens of seconds.
2.
Non-decreasing luminosity: Under the assumption that only SM particles are emitted and no additional degrees of freedom are involved, the luminosity increases steadily until the final moment of evaporation.
3.
Locally-constrained signal: Given the sensitivities of current instruments, PBH bursts are expected to be detected within a few parsecs of Earth.
4.
Universal iontrinsic spectrum: The intrinsic emission spectrum is characteristic of the evaporation process, implying that different PBH bursts should appear spectrally identical.
Several instruments have set upper limits for the PBH burst-rate density, . In the case of Cherenkov telescopes such as H.E.S.S. and VERITAS, the upper limits are pc-3 yr-1 with confidence level, and pc-3 yr-1 with confidence level respectively (Aharonian et al., 2023a; Archambault, 2017). In the case of wide-FoV WCDs, HAWC has established an upper limit pc-3 yr-1 with confidence level using data of 3 years with events located at pc (Albert et al., 2020). Most recently, prospective observations of LHAASO have set more constraining upper limits with pc-3 yr-1 for 5 years of data with confidence level, considering bursts at distances pc from the Earth (Yang et al., 2024).
SWGO’s approach to PBH burst detection (or the setting of upper limits) will be based on previous gamma-ray detector strategies. This will consist of counting the number of photons, , from a PBH burst with an evaporation time inside a volume of radius , and can be expressed as:
where is the true (reconstructed) energy, is the effective area, is the resolution, and is the dead time of the detector (given SWGO’s high duty cycle, will be negligible). There will be particular focus on detecting very short bursts in order to capture emission from the TeV energy band and to minimize dependence on the zenith angle of the source. Access to the southern sky plus the improved sensitivity of SWGO compared to other WCD facilities present a major opportunity to search for signatures of PBH evaporation and to further constrain existing upper limits (López-Coto et al., 2021). Although this is a serendipitous phenomenon, its detection would have a profound implication for fundamental physics through its interplay in cosmology (evidence of non-stellar BH formation), high-energy physics (probing BSM physics), and gravitation (testing Hawking radiation as a key quantum gravitational phenomenon).
7.3 Axion-Like Particles (ALPs)
ALPs are predicted in many extensions of the Standard Model and another promising avenue for exploring BSM physics via gamma-ray observations. ALPs are hypothetical particles that arise in BSM theories (Jaeckel & Ringwald, 2010), related to the axion (Peccei & Quinn, 1977), and are considered viable DM candidates (Arias et al., 2012). They are extremely light particles that interact weakly with matter, rendering their direct detection highly challenging. Although still hypothetical, ALPs may provide valuable insights into the Universe and contribute to resolving outstanding questions such as the nature of DM.
The relation between ALPs and photons is described by the Lagrangian:
On the right-hand side of equality, the first term, , can be expressed as (Meyer et al., 2013):
It represents the coupling between photons and ALPs, where and are the electromagnetic field tensor and its dual, respectively, is the ALP field, and is the ALPs-photons coupling factor. The second term, is the Euler-Heisenberg correction Lagrangian, which describes photon propagation in the presence of a strong electromagnetic field (Meyer et al., 2013). The third term, , can be expressed as:
which depends on the kinetic energy and the ALP mass, . Therefore, each ALP candidate can be uniquely defined by the parameters .
Extragalactic sources are typically not observed at very high energies on Earth because their spectra are attenuated by the EBL. However, if the gamma rays convert to ALPs while traversing magnetic fields in the intergalactic medium, and subsequently reconvert into gamma rays when they interact with the Milky Way’s magnetic field, they can survive EBL absorption and be detected on Earth (Horns et al., 2012; Meyer et al., 2013). The probability of conversion, , is described as:
where is the ALP critical energy, is the transversal component of the magnetic field, and is the distance traveled within the ALP-interaction region. Thus, the photon flux observed at Earth results from the convolution of the emitted flux at the source, and the ALP-photon conversion probability following:
If ALPs exist, their conversion into photons could allow for the detection of gamma rays that would otherwise be absorbed. A potential observational signature of ALPs may involve the detection of photons with energy 30 TeV coming from a hard-spectrum extragalactic source, or 10 TeV from soft-spectrum galactic sources. These high-energy regimes are particularly suited for indirect searches with SWGO whose energy range is expected to go beyond 100 TeV.
The existence of ALPs is still an open question in current physics; their indirect detection from astrophysical sources would allow us to test various theoretical models and improve our understanding of the systematic uncertainties associated with the EBL models (Franceschini et al., 2008; Domínguez & et al., 2011; Finke et al., 2010) and, by extension, the intrinsic energy spectrum.
A particularly relevant source for study is the AGN Centaurus A (Cen A) due to its intense emission of very-high-energy gamma rays, its proximity () to Earth, and the astrophysical conditions favorable for photon-to-ALP conversion in the intergalactic magnetic field (H. E. S. S. Collaboration, 2020, 2018a). The emission of this point-like extragalactic source has been modeled as a power-law spectrum with index, and normalization, at an energy of 1 TeV (H. E. S. S. Collaboration, 2019). Photons up to TeV are expected, even accounting for EBL absorption. With its wide energy coverage extending beyond 100 TeV, SWGO offers the opportunity to probe emissions from Cen A, potentially providing indirect evidence for photon-ALP conversion. Figure 26, left panel, shows the expected spectrum, assuming an emitted power-law, of Cen A with the full SWGO array for 5 and 10 years of observation. Figure 26, right panel, shows the exclusion regions in the phase-space, computed using these two observation times, and in comparison with exclusion regions from different experiments. For these simulations, the EBL model from Franceschini et al. (2008) was assumed. As the Figure shows, SWGO will probe new ALP parameter space with such a Cen A observation. In addition, its wide FoV provides the potential for finding new targets for such ALP searches.
7.4 Testing Lorentz Invariance
Very high energy (VHE) photons serve as pivotal probes in testing fundamental physics principles, such as Lorentz symmetry. Examining the limits of any symmetry’s validity is essential for theoretical advancements and experimental physics. These investigations are motivated by the hypothesis that Lorentz invariance violation (LIV) may arise from BSM theories, including quantum gravity and string theories (et al. (2022); Martínez-Huerta et al. (2020); Kostelecky & Russell (2011) and references therein).
Previous studies indicate that LIV may introduce modifications in the propagation of high-energy photons, including changes in their velocity, interaction thresholds, and arrival times. These effects can be modeled by modifying the photon dispersion relation, which takes the form:
where and are the energy and momentum of the photon, respectively, and denotes the scale of Lorentz invariance violation at order . These scales indicate the energies at which corrections of the corresponding order become relevant.
The sign ambiguity in Equation 14 gives rise to two distinct scenarios: superluminal () and subluminal (). Phenomenological consequences of such modifications are typically studied on a term-by-term basis for specific orders, with a focus on the linear () and quadratic () cases, as these yield the most significant effects and are more readily constrained by observations. The observable consequences of such modifications include, among others, anomalies in particle interaction thresholds, photon decay, photon splitting, birefringence effects and time delays (et al., 2022). For instance, in the linear case, birefringence (where photons of opposite polarization states propagate at different speeds) can place a lower bound on the energy scale at which new physics may emerge, , many orders of magnitude above the Planck scale, eV (Wei, 2019; Gotz et al., 2014).
In the sub-luminal scenario of Equation 14, attention is given to anomalies in particle interaction thresholds (Scully & Stecker, 2009; Stecker & Scully, 2009; Galaverni & Sigl, 2008). These effects alter the kinematics of pair production, , the key process governing the attenuation of VHE gamma rays, where the second photon originates from low-energy background media, such as the CMB or the EBL. Specifically, the threshold energy for this pair-production process is modified. For head-on collisions, the threshold energy, , of the low-energy photon is given by:
where the second term on the right-hand side, introduced by LIV, results in the appearance of a minimum in . In contrast to the Special Relativity (SR) scenario, where the energy threshold decreases monotonically with increasing gamma-ray energy, the presence of a minimum in leads to greater transparency of the background medium beyond a certain critical gamma-ray energy (Jacob & Piran, 2008). The critical energy of gamma rays corresponding to this minimum is determined by differentiating Equation 15. For the linear case, the critical energy is
Given the strong constraints arising from birefringence effects, the quadratic case becomes particularly interesting as a phenomenological window to explore. The interest is further motivated by a recent first-principles calculation of the cross-section for the pair-production process (Carmona et al., 2024). This result contrasts with previous approaches, which relied on approximations based on the cross-section in the SR scenario. Such approximations have been shown to overestimate LIV-induced transparency of the background, thereby underestimating the scale of LIV. In this case, the critical energy can be expressed as:
Thus, observations at energies PeV with SWGO, we will be probing LIV scales nearly four orders of magnitude larger than current constraints. Current bounds on from anomalous transparency studies, derived from single-source analyses, stand at approximately eV (), based on observations of gamma rays up to 20 TeV from the Mrk 501 flare (Abdalla et al., 2019).
In super-luminal scenarios, Equation 14 permits photon decay into electron-positron pairs above a certain energy threshold, leading to rapid decay of high-energy photons (Martínez-Huerta & Pérez-Lorenzana, 2017). This effect would result in a suppression of photon emission from sources exceeding this energy threshold, which would manifest as a hard cutoff on the high-energy tail of spectral sources (Albert et al., 2020). As a result, the non-observation of such a cutoff in measurements of astrophysical VHE photons constrains the energy scale at which photon decay becomes kinematically allowed, based on the observation of photons with energies up to . This allows us to refine the bounds on the LIV energy scale, given by:
In this context, observing photons with energies ranging from 1 to 5 PeV by SWGO, may substantially enhance constraints on LIV derived from Southern Hemisphere observations. The absence of decay signatures can be used to establish limits on LIV, potentially to or beyond the Planck scale, , depending on the decay model (linear or quadratic). For a linear modification, the resulting super-luminal LIV bounds may range from eV to eV. For a quadratic modification, using similar calculations, they range from eV to eV. Current stringent super-luminal limits for are on the order of eV (Li & Ma, 2022).
Additionally, photon splitting processes, a phenomenon where high-energy photons split into multiple lower-energy photons under LIV scenarios with , may also offer additional constraints on LIV (Gelmini et al., 2005; Satunin, 2019). Observations of sources like RX J1713.7-3946 without any signatures of photon splitting at energies eV can further limit the LIV scale, as predicted by:
The calculated LIV limits for photon splitting at eV and eV are eV and eV, respectively, providing further empirical support for refining the bounds on LIV.
A comprehensive exploration of both sub-luminal and super-luminal regimes is essential to systematically test the full range of predictions arising from Lorentz-violating extensions. Addressing both cases independently enhances the interpretative power of observational constraints and ensures sensitivity to the diverse phenomenology predicted by different theoretical frameworks. Furthermore, as explored in Section 6, gamma-ray observations of transients may also serve as critical probes for LIV scenarios and ALP-induced modulations.
8 Cosmic-Ray (CR) Measurements
In addition to a vast range of gamma-ray topics, SWGO will act as a unique tool for studying charged cosmic rays in the TeV-PeV regime. Its large effective area and advanced muon counting capability, combined with its privileged location in the Southern Hemisphere, will allow the pursuit of relevant open questions in TeV-PeV cosmic-ray astrophysics. The main topics of interest are described in detail below.
8.1 Spectrum and Composition of CRs around the Knee
The mechanism behind the acceleration of CRs and the physical processes that affect their propagation through space are subjects of intense debate. An indirect approach to investigating these is to study the CR energy spectrum and mass composition. In the energy region around and below the knee (1 TeV 10 PeV), where SWGO will be sensitive, there have recently been several experimental efforts to study the energy spectrum the bulk of cosmic rays and the relative abundances elemental mass groups of this radiation. These aim to fill the gap between direct and indirect CR data and to refine the measurements around the knee to investigate the properties of galactic CRs. Between TeV and PeV, measurements on the all-particle spectrum of CRs from the NUCLEON satellite (Atkin et al., 2017; Grebenyuk et al., 2019) and the HAWC observatory (Alfaro et al., 2017, 2025) have established a first bridge with high statistics and large precision between the TeV CR balloon- and space-borne experiments and the PeV air-shower observations and have showed the existence of a softening at tens of TeV. The last analysis of HAWC has located this structure at TeV (Alfaro et al., 2025). More recently, LHAASO has measured the total spectrum around the knee, from TeV to PeV, with a precision which have not been achieved so far by other experiments (The LHAASO Collaboration et al., 2025). The position of the knee was found at PeV. LHAASO spectrum is in agreement with HAWC data between TeV and 1 PeV and shows no features below the knee.
Measurements of the elemental mass groups of CRs in the TeV energy region have also recently discovered several breaks in their energy spectra. NUCLEON (Grebenyuk et al., 2019), DAMPE (An et al., 2019; Alemanno et al., 2021) and CALET (Adriani et al., 2019, 2023) have observed the existence of cutoffs in the CR spectra of light elements: close to TeV for protons (or H nuclei) and around TeV for He primaries. DAMPE Alemanno et al. (2024) and HAWC Albert et al. (2022) have also observed a softening in the spectrum of HHe nuclei at . Preliminary results from HAWC point towards the TeV softening of the all-particle CR spectrum is due to the cutoffs in the H and He components of CRs and an increment in the contribution of heavy primaries Arteaga-Velázquez et al. (2024). Hardenings at around in the spectra of H and He nuclei have been also reported by HAWC Arteaga-Velázquez et al. (2024). GRAPES-3 has confirmed such feature for the spectrum of protons at TeV Varsi et al. (2024). LHAASO has provided recent data on the proton spectrum with the largest precision to date within the energy range and has detected a hardening close to TeV (Cao et al., 2025b). It should be pointed out that proton flux measured by LHAASO is larger than that measured by GRAPES-III and HAWC between TeV and PeV. Regarding heavier CR primaries, NUCLEON results have hinted at the presence of individual cutoffs in the elemental spectra of heavy nuclei at rigidities (Atkin et al., 2017). HAWC has also pointed out a softening in the spectrum of the heavy (C-Fe) component of CRs at hundreds of TeV Arteaga-Velázquez et al. (2024).
At energies around the knee, measurements of the air-shower experiments KASCADE (Antoni et al., 2005; Apel et al., 2009, 2013), IceTop / IceCube (Aartsen et al., 2019) and MAKET-ANI (Chilingarian et al., 2004, 2007) indicate that this softening is produced by cutoffs in the spectra of light primaries (H and He nuclei). Also, the latest results of LHAASO on the mean logarithmic mass supports the picture that the knee is dominated by light primaries (The LHAASO Collaboration et al., 2025), and the measurements on the energy spectrum of protons confirm the existence of a cutoff at energies around the knne ( PeV) (Cao et al., 2025b). However, results from the ARGO-YBJ (Bartoli et al., 2015; Montini et al., 2016) and Tibet-ASgamma (Amenomori et al., 2019) experiments show a cutoff in the spectrum of the H+He mass group lies at around 400-700 TeV. More accurate measurements of the elemental spectra over the full energy range are needed to determine whether this is an additional feature or whether this is in tension with the other experimental results of the knee being due to a softening of light CRs. KASCADE-Grande data have also shown individual cutoffs in the spectra of the CNO and Fe components. Interestingly, within systematic and statistical errors, the positions of the breaks detected by KASCADE-Grande in the elemental groups of CRs seem to depend on the charge of the primary particle (Antoni et al., 2005; Apel et al., 2009, 2013).
The results of KASCADE-Grande on the knee-like features in the spectra of the mass groups point to a rigidity dependence consistent with a Peters cycle (Peters, 1961), as the maximum acceleration energy of a source population is expected to be proportional to the nuclear charge. However, also propagation effects on the cosmic-ray spectra are expected to depend on the rigidity, and the knee would then indicate the point where galactic PeV CRs start to escape from their magnetic confinement at the sources (Hörandel, 2004; Cristofari et al., 2020; Vink, 2022) and/or in the Milky Way (Candia et al., 2002; Hörandel, 2004; Giacinti & Semikoz, 2023). The origin of the TeV knee-like feature in the all-particle spectrum of CRs may have a similar origin according to NUCLEON results. If confirmed, this scenario would predict charge-dependent cutoffs in the TeV CR spectra of the heavier nuclei, too. For the CNO and Fe mass groups, the cutoffs would lie between and TeV. The existence of these features could indicate the presence of an individual nearby TeV CR source inside our galaxy (Liu et al., 2019) or a complete population of such a kind of CR sources (Yue et al., 2020). It could also point out an unknown mechanism affecting the propagation of TeV CR in the Milky Way (Malkov & Moskalenko, 2022). The hardenings detected at around TeV by HAWC Arteaga-Velázquez et al. (2024) and Grapes-3 Varsi et al. (2024) in the spectra of the light CR nuclei may indicate the point where the contribution from these phenomena is surpassed by the contribution from the background of PeV CR sources.
In order to test different hypotheses about the origin of CRs at TeV and PeV energies, the shape of the energy spectra and the location of the knee-like features for the elemental mass groups of CRs must be measured with high precision. SWGO can contribute to this task by measuring the elemental spectra in the TeV energy regime for H, He, CNO and Fe primary nuclei. The H and He TeV cutoffs need confirmation with large statistics and high precision, as well as the CNO and Fe TeV softenings, which have just been hinted at by some experiments. SWGO data would also be important in investigating the transition from the TeV to the PeV energy regime in the spectra of CRs by measuring the TeV hardenings in the spectra of CRs with high precision. In addition, SWGO can contribute to measuring the precise location of the p, He, and CNO knee-like features at PeV energies to test models about the production and propagation of galactic CRs. LHAASO results on the energy spectrum of protons and on the abundance of light primaries around the knee need to be confirmed, as well as the HAWC results on the origin of the TeV cutoff at the total CR spectrum. It is important to point out that SWGO measurements in the TeV energy range will also be important to cross-check the energy scales of current direct and indirect experiments.
SWGO will offer insight into the above questions by: (a) providing a large effective area and high duty cycle for the detection of CRs in the TeV-PeV energy regime, and (b) taking advantage of improved muon-counting capabilities compared to current instruments that are sensitive to the same energy range e.g., HAWC. An estimate of the proton effective area achievable with SWGO is shown in Figure 27. For this estimation, we apply conservative selection criteria, requiring that the shower core falls within the array boundaries, a minimum of 25 triggered (or hit) tanks, and a zenith angle between and . We define the energy threshold as the energy for which the effective area reaches 10% of its saturated value. The effective area for the SWGO (SWGO-A) configuration reaches , with an energy threshold, GeV. This enables SWGO to achieve the high event statistics necessary for precise studies in the TeV–PeV energy range.
In addition, enhanced muon-counting capabilities will improve the determination of the primary CR composition. Generally, identifying the species of a CR primary using WCDs is challenging, as the longitudinal profile of the shower is not measured directly. However, the dual-layer tank design makes the larger top layer act as a calorimeter for the electromagnetic component of the shower while allowing the muons to reach the bottom layer, making it a crucial tool for identifying and, ideally, counting them. The relation between the total electromagnetic energy, , and the total number of muons, , of an air shower may be used as a strong separator between different CR species. See Section 8.3 for more details regarding SWGO’s study of the muon lateral distributions.
The possibility of separating the composition into four main groups, H-like, He-like, N-like, and Fe-like, is desirable. To that end, different methods for estimating and have been explored in the context of SWGO, with the most promising one being a template-based search (Lang et al., 2023). Template-based reconstructions have already proven to be accurate for gamma-ray events detected with WCDs (Joshi et al., 2019) and are one of the methods considered for core, energy, and direction reconstruction of gamma rays in SWGO. For CR studies, we build templates as probability density functions (PDFs). Monte Carlo simulated events are binned in true and and used to fill histograms of measured charge (in photoelectrons or p.e.) vs. distance to the core of the shower in the shower plane. Each of the four representative primary particle groups is treated individually for building the templates. Thereafter, the charge distribution of an event is compared to these PDFs, resulting in a likelihood of such an event having a given and/or . Preliminary implementations of the method have shown to reach a resolution in the number of muons of the order of 10-20% for He-like, N-like, and Fe-like, and of the order of 30% for protons Lang et al. (2023). The combined and likelihoods are then used to discriminate between the most likely primary particle group and background. While more realistic conditions, such as uncertainties in core reconstruction, still need to be accounted for, current estimates indicate that using a dual-layer tank significantly improves primary CR separation, achieving a purity of nearly 90% for protons and at least 95% for heavier primaries (Lang et al., 2023). This will allow SWGO to measure the energy spectra for different mass groups, helping to resolve existing tensions in CR composition measurements around the knee and providing crucial insights into the origin of PeV CRs.
8.2 Anisotropy
Due to their electric charge, CRs are deflected by magnetic fields, resulting in strongly diffusive propagation within our galaxy. This process masks out information about their original acceleration sites. Therefore, the distribution of arrival directions of CRs (on Earth) is highly isotropic. Nevertheless, ground-based observatories covering large areas have recently reached sufficiently large exposures to resolve small but significant anisotropic structures. These anisotropic structures are typically characterized using the angular power spectrum derived from a spherical harmonic decomposition of the arrival direction distribution. For the TeV-PeV range, the most prominent deviation from isotropy comes in the form of a significant but small dipole, with an amplitude of the order (Abbasi et al., 2025).
The study of large-scale anisotropies can lead to strong biases if only partial FoVs are considered. Ground-based instruments have access only to a fraction of the sky, and complementary measurements from different experiments are desirable for achieving full-sky coverage. This was first done in the TeV region by HAWC and IceCube (Abeysekara et al., 2019). However, combining data remains challenging due to differences in detection techniques as well as sensitivities to energy and composition across experiments. Any attempt to align the systematic uncertainties is further complicated by the small region of overlap ( for the HAWC and IceCube measurements) in which the direction reconstruction quality is the poorest for both experiments. SWGO will contribute significantly to studying the full-sky dipole anisotropy by providing large statistics and high-quality data, employing a detection mechanism similar to that of HAWC and a larger overlap with its FoV. For all considered primaries, even a simple plane fit reconstruction for the arrival directions can help achieve angular resolutions , enough to explore mid-scale anisotropies, for which the size of structures is of the order of tens of degrees.
The amplitude and phase of the measured dipole component also present an interesting evolution with energy, as already observed by IceCube (Abbasi et al., 2025). A “swing” in the dipole component direction is observed between 10 TeV and 1 PeV – its phase in right ascension gradually changes from pointing away from the GC for TeV to pointing roughly close to the GC for TeV. Correspondingly, a drop in amplitude is seen between 25 TeV and 100 TeV followed by a steep increase up to a few tens of PeV. Moreover, increasing multipole contributions become evident at higher energies (Abbasi et al., 2025).
Although the origin of the observed dipolar anisotropy remains an open question, its energy-dependent evolution and its decomposition by mass group may hold the key to understanding its nature. If the anisotropy arises from propagation effects, one would expect this evolution (or “swing”) to appear at different energies for different CR nuclei. Furthermore, magnetic field effects are modulated by particle rigidities, implying that a rigidity-dependent behavior (instead of an energy-dependent one) is foreseen. Therefore, a reasonable event-by-event species classification is desired. As described in Section 8.1, a template-based method has demonstrated effective separation of CR primaries into four mass groups. A confusion matrix can also be constructed to estimate the probability of a CR primary being misclassified into an incorrect mass group, for a given energy and species. This enables the application of unfolding procedures to correct for classification errors. As a result, its enhanced muon-counting capabilities will allow SWGO to perform pioneering measurements of composition-dependent anisotropy in the TeV-PeV energy range, offering new insights into the origin of the galactic CR dipole.
In literature, large-scale anisotropy (e.g., the dipole component) is usually associated with diffusive propagation of CRs through interstellar magnetic turbulence (Blasi & Amato, 2012; Ahlers, 2016; Qiao et al., 2022; Li et al., 2024) and may, therefore, provide crucial insights into the properties of this turbulence (Giacinti & Sigl, 2012; Giacinti & Kirk, 2017; Giacinti & Reville, 2022; Bian et al., 2024). However, the observed CR anisotropy exhibits a complex angular scale structure, typically described using spherical harmonic functions. This mathematical description enables investigation into the origin of these observations, including non-diffusive propagation effects occurring within the mean scattering length, correlations among particles traversing the same turbulent magnetic fields, and the effects of turbulent convection (Ahlers, 2014; Ahlers & Mertsch, 2015, 2017; Mertsch & Ahlers, 2019; Génolini & Ahlers, 2021; Kuhlen et al., 2022; Zhang & Liu, 2024). By offering an unbiased perspective on the distribution of power across various angular scales, this approach is essential for testing and distinguishing between various models that explain the observed medium- and small-scale anisotropies. SWGO will further enhance studies on the properties of interstellar turbulence by leveraging its detection redundancy and providing high-quality CR data, thereby advancing our understanding of CR propagation.
8.3 Muon Puzzle and Hadronic Interaction Models
The inference of the mass composition of air showers is deeply intertwined with our understanding of the hadronic interactions that govern their development. To accurately model the shower evolution, hadronic interaction models must describe the full phase-space of multiparticle production. This relies on phenomenological parameterizations that are calibrated to available accelerator data and extrapolated to higher energies and kinematic regions that lie beyond current experimental reach.
Despite extensive efforts, state-of-the-art hadronic interaction models, tuned to data from the Large Hadron Collider, fail to accurately predict the muon content of air showers at ultra-high energies (Aab et al., 2015, 2016; Abdul Halim et al., 2024). This discrepancy, commonly referred to as the Muon Puzzle, has been observed across multiple experiments, suggesting that the divergence between data and simulations may emerge as early as at PeV energies (Arteaga Velazquez, 2023).
Measurements of the relative fluctuations in the muon number from the Pierre Auger Observatory (Aab et al., 2021) align with mass composition expectations derived from the analysis of the depth of the shower maximum in the longitudinal profile. Combined with the findings of Cazon et al. (2018), this evidence strongly suggests that the origin of the Muon Puzzle may lie in subtle deviations across numerous low-energy hadronic interactions, rather than in a substantial modification at the highest energies. Additionally, muon distributions at low energies have been found to depend on the choice of both – low- and high-energy hadronic interaction models (Pastor-Gutiérrez et al., 2021).
For all these reasons, SWGO must validate the modeling of shower physics and accurately measure the muon component of EAS at ground level. In fact, with its capability to identify EAS muons – necessary to perform composition studies and gamma/hadron separation – and high fill factor of detectors, SWGO is particularly well-suited for this task in the energy range from a few hundred GeV to several tens of PeV. It is also worth noting that the transition between low- and high-energy hadronic interaction models occurs at (100 GeV), further enhancing the relevance of this energy regime.
In the following, we describe one possible strategy that SWGO can employ to discriminate between high-energy hadronic interaction models. As described in Section 8.1, the template-based method used for the energy reconstruction and composition separation relies on the muon lateral distribution functions (LDFs) derived from simulations and is, therefore, highly sensitive to variations in hadronic interaction models. However, the correlation between measurements from the bottom and top layers of the detector offers a nearly independent estimate of the number of muons in each individual tank. Since this method does not depend on the overall shower distribution, it is less sensitive to uncertainties in hadronic interaction models. Differences in the muon energy spectrum and the punch-through effect are expected across models, however they represent second-order effects.
Preliminary studies (Kunwar et al., 2023) have demonstrated the feasibility of using the bottom layer as a muon tagger, allowing for the identification of tanks where one or more muons contribute to the signal. This capability is important for gamma/hadron separation, as the expected number of muons in gamma-ray air showers is very low. In the case of CRs, however, the number of muons increases substantially with energy, reaching muons at the detector level for Fe nuclei at TeV energies. In this regime, precise muon counting, rather than mere tagging, becomes essential.
An initial assessment of SWGO’s ability to estimate the muon LDF from individual tank measurements, largely independent of hadronic interaction models, can be obtained by analyzing the integrated charge in each layer of every tank. Figure 28 presents the median integrated charge from the top and bottom layers as a function of the distance to the shower core in the shower plane for several proton-induced events within an example energy range of 980 TeV up to 1000 TeV, simulated using two different hadronic interaction models: QGSJET-II.04 (Ostapchenko, 2011) and EPOS-LHC (Pierog et al., 2015). The predicted LDFs for the measured charges exhibit notable differences between the two models, highlighting SWGO’s potential to place strong constraints on hadronic interaction models and contribute to a more model-independent investigation of the muon LDF and the Muon Puzzle.
The impact of uncertainties in energy reconstruction, arising from assumptions about hadronic interaction models in such a simplified analysis, remains to be evaluated. Nevertheless, more sophisticated and robust approaches to individual tank muon counting can be developed using machine learning techniques. For instance, deep neural networks (DNNs) have demonstrated improvements in gamma-ray reconstruction for detectors like SWGO Assis et al. (2022); de Almeida et al. (2023); Conceição et al. (2024); Glombitza et al. (2025b), suggesting their potential for enhancing muon detection as well.
8.4 Heliosphere and Solar Physics
The Sun has emerged in recent years as a new source of interest in gamma-ray astronomy, with a puzzling very-high-energy signal observed by Fermi-LAT and HAWC in the GeV-TeV range (Abdo et al., 2011; Albert et al., 2023a). Nominal models of CR interactions with the Sun’s atmosphere are unable to explain the hard spectrum and morphology of these solar gamma rays. High-statistics gamma-ray measurements of the Sun and a potential resolution of the solar gamma-ray puzzle would be a significant step forward in our understanding of local CR propagation in the Sun’s dynamic environment, its magnetic fields, as well as in searches for dark matter and new physics (Bell et al., 2021). The Sun being a bright, moving source can only be efficiently probed using an all-sky survey instrument that is capable of day-time operations, effectively ruling out observations from even the most sensitive IACTs. SWGO, with its improved sensitivity over HAWC, is perfectly suited for a precise measurement of the gamma-ray spectrum of the Sun during future solar cycles and will enable the search for additional temporal features in the emission. Figure 29 shows the existing measurements of the Sun in gamma rays along with the projected sensitivity of SWGO.
9 Multi-Messenger and Multi-Wavelength (MM/MWL) Program
SWGO will offer unprecedented capabilities for high-energy time-domain astronomy from the Southern Hemisphere. With its wide FoV, high duty cycle, and continuous monitoring strategy, SWGO is uniquely positioned to serve as a cornerstone facility for MM/MWL astrophysics. It will offer crucial Southern sky coverage, complementing Northern facilities such as LHAASO, and will play a vital role in the global network of observatories probing the transient Universe. SWGO’s design and operational strategy will address several major challenges in time-domain and MM/MWL astrophysics in the following ways:
High Duty Cycle: The observatory is expected to operate with a duty cycle exceeding 95%, enabling near-continuous observation of the sky. Unlike IACTs, which are limited by weather, daylight, and moonlight, SWGO will operate both day and night, regardless of atmospheric conditions. This makes it ideally suited for detecting unpredictable, short-lived, or recurrent astrophysical events.
Large Instantaneous FoV: With an FoV of approximately 2 steradians, SWGO will monitor a substantial portion of the sky at any given moment. This will increase the likelihood of detecting rare and serendipitous transients across a broad range of spatial and temporal scales. It also enhances SWGO’s potential to discover new classes of variable and transient gamma-ray sources.
All-Sky Monitoring without Repointing: SWGO’s design permits it to continuously observe the entire overhead sky without the need for repointing or scheduling (as opposed to IACTs). This ensures the availabilty of pre-alert data for retrospective analysis (following external alerts), making SWGO highly effective for identifying not only precursor activity but also significantly delayed high-energy counterparts.
Real-Time Data Analysis: SWGO will implement real-time pipelines for both externally triggered follow-ups and blind searches for transient events. These systema will facilitate rapid detection of gamma-ray counterparts to high-energy neutrinos, GWs and FRBs, as well as unexpected phenomena such as magnetar flares, galactic transients, and unidentified bursting sources. The ability to respond within seconds to minutes of an external alert will place SWGO at the forefront of rapid-response high-energy astrophysics.
SWGO’s effectiveness in time-domain and MM/MWL astrophysics will rely on how it shares data with the global astrophysics community. As a modern observatory, it must be fully embedded within the global MM ecosystem. With this in mind, SWGO will adopt a transparent and community-oriented data policy, designed to maximize accessibility and scientific return:
Real-Time Alert Dissemination: Based on the real-time data analysis, the observatory will issue alerts containing key event parameters, such as time, localization (with uncertainty), false alarm rate (FAR), energy range, and flux estimates (or upper limits), using standardized formats like the International Virtual Observatory Alliance (IVOA) VoEvent schema. This will ensure a rapid interpretation of SWGO data and responses from other observatories and automated alert brokers.
Archival Data and Source Catalogs: On longer timescales (typically within one year), SWGO will release comprehensive datasets from offline analyses, including transient light curves and catalogs of steady and variable sources. These products will be made available through FAIR-compliant platforms, integrated into established databases such as VizieR, HEASARC, and TeVCat (Wakely & Horan, 2008), to ensure interoperability and long-term discoverability.
9.1 Synergies with the Global MM/MWL Ecosystem
The astrophysical transient landscape is rapidly evolving, with a growing number of facilities generating real-time alerts across the electromagnetic spectrum and beyond. SWGO is designed to contribute to and benefit from this ecosystem. Its engagement with major alert networks will ensure timely responses to external triggers and allow the broader community to act on SWGO’s findings. Key platforms in this collaborative framework include:
Astronomer’s Telegrams (ATels; Rutledge (1998)): ATels are human-curated reports on variable sources, such as AGN flares or X-ray binaries, which will be monitored by SWGO for VHE gamma-ray activity on timescales of days or longer. SWGO will contribute its own findings back to the community via ATels, ensuring rapid dissemination of significant detections.
General Coordinates Network666https://gcn.nasa.gov/ (GCN): SWGO will respond to real-time GCN alerts from instruments detecting GRBs, high-energy neutrinos (e.g. IceCube), and gravitational-wave events (e.g. LIGO-Virgo-KAGRA). Its ability to instantaneously monitor relevant sky regions will enable prompt detections of potential counterparts (including refined localizations of poorly localized events) or the setting of upper limits.
Astrophysical Multimessenger Observatory Network (AMON; Smith et al. (2013)): Building on experience with HAWC, SWGO will contribute real-time data streams of sub-threshold event clusters and selected high-energy photon candidates. This input will help identify statistically significant coincidences across different wavelengths and messengers, facilitating early detection of flaring states or entirely new source classes.
Transient Name Server777https://www.wis-tns.org (TNS): SWGO will monitor reports from the TNS to follow up on optical transients such as supernovae, TDEs and FRBs, particularly those associated with non-thermal or rebrightening activity.
Vera Rubin/LSST Brokers888https://www.lsst.org/scientists/alert-brokers: With the Rubin Observatory expected to detect millions of transients per night, real-time classification and follow-up will be critical. SWGO will be the only ground-based instrument that will cover the same sky simultaneously with LSST, therefore, ideally placed to search for gamma-ray counterparts to LSST transients over a wise range of timescales from minutes to weeks and months. As with AMON (see above), SWGO will employ dedicated pipelines to correlate its sub-threshold stream of gamma-ray events with LSST-detected transients.
Astro-COincidence LIBrary for Real-time Inquiry (Astro-COLIBRI; Reichherzer et al. (2021)): This centralized coordination platform for transient astrophysical events, developed within the MM community, will serve as a primary interface for SWGO’s burst advocates and shift personnel. It will allow real-time filtering, visualization, and contextualization of alerts across all relevant messengers and wavelengths.
In addition to its scientific role, SWGO will foster broader engagement within the astronomical community and beyond. It will leverage the vibrant local amateur astronomy community in the regions around San Pedro de Atacama for long-term photometric and spectroscopic monitoring of gamma-ray sources. These collaborations, coordinated through platforms like Astro-COLIBRI, will expand the observational reach of SWGO and also provide valuable training and outreach opportunities. Through partnerships with initiatives like the United Nations’ Open Universe999https://openuniverse.cbpf.br program, SWGO will contribute to capacity building in data science and astrophysics across Latin America and the Global South, ensuring that its impact is as inclusive as it is scientifically significant.
9.2 High-Energy Neutrinos
The combination of gamma-rays with high-energy neutrinos is extremely powerful for the identification of high-energy CR accelerators, given the likely association of neutrino signals with hadron acceleration, and the larger statistics and precision possible in the gamma-ray. Extensive experimental attempts in recent decades have culminated in the finding of a diffuse neutrino flux by the IceCube Collaboration (Aartsen et al., 2013; Abbasi et al., 2021; Abbasi et al., 2022) as well as the first neutrino sources, including the flaring blazar TXS 0506+056 (IceCube Collaboration et al., 2018), the Seyfert galaxy NGC 1068 (IceCube Collaboration et al., 2022), and the Galactic Plane (Abbasi et al., 2023). Despite the emergence of the first source classes, the origin of the IceCube diffuse flux remains a mystery.
Located in the Southern Hemisphere, SWGO is uniquely placed to follow up IceCube’s Southern-sky tracks and Earth-skimming events, as well as KM3NeT’s Earth-crossing events, and will provide a map of the TeV-PeV photon sky which will be invaluable for understanding the galactic neutrino emission in the same energy range.
SWGO’s unprecedented sensitivity to both point sources and diffuse emission from the Galactic Plane in the southern sky will enhance the galactic program of neutrino experiments by providing valuable templates and source models (Fang & Halzen, 2024). It will also observe promising neutrino source candidates such as nearby blazars, tidal disruption events, and star-forming galaxies. Additionally, SWGO may help constrain neutrino production in GRBs by detecting more nearby, luminous bursts.
9.3 Gravitational Wave (GW) Sources
Observations by the LIGO, Virgo, and KAGRA interferometers have established the field of gravitational-wave (GW) astronomy, marked by several key milestones: the first direct GW detection (GW150914), the first binary neutron star (BNS) merger with an electromagnetic counterpart (GW170817), an intermediate-mass black hole merger (GW190521), an unusual mass-ratio system (GW190814), and the first neutron star–black hole (NS–BH) merger (GW200115).
Despite ongoing advances in GW detectors, events with electromagnetic counterparts, such as BNS mergres and NS–BH mergers, will remain confined to the local universe (), a redshift range that SWGO can probe despite EBL absorption. With its high duty cycle and continuous monitoring abilities, SWGO will be well-positioned to follow up GW events in the Southern sky. As localization regions for GW detections are expected to remain large (110–180 deg2), SWGO’s wide FoV makes it one of the few gamma-ray instruments capable of efficiently covering them. Moreover, SWGO’s real-time alert system could enable prompt follow-up by facilities like CTAO, ensuring broad multi-wavelength coverage.
9.4 Combination with Cosmic Rays Measurements
Many SWGO observations relate directly to cosmic particle acceleration as discussed elsewhere in this document. Direct connections to the cosmic ray observatories include SWGO’s capability to observe CR anisotropy is discussed in Section 8.2, but also complementarity to the ultra-high energy (UHE) CR observatories.
The higher-end of the SWGO energy range is limited to the GZK horizon by EBL absorption and can hence complement neutrino studies in probing the accelerators responsible for the locally measured UHECRs (see for example Albert et al. (2022)).
The Southern Hemisphere location makes SWGO ideally suited to probe the most significant deviation from isotropy in the UHECR sky: an excess likely associated to the very nearby active galaxy Cen A. The combination of sensitivity to many degree scale emission and in to the PeV range is extremely promising in this regard.
The combination of SWGO mass-resolved anisotropy measurements with a wide range of CR observatories in the TeV–PeV range, will be extremely powerful as a probe of local CR transport as well as CR origin.
10 Outreach & Communications
In addition to the strong scientific objectives that constitute the pillars of the SWGO project, the collaboration has devoted particular attention to the development of a comprehensive outreach and education program. Through our official website www.swgo.org and YouTube channel, the Collaboration actively disseminates updates and information about ongoing activities in a way that is accessible to the general public. Following the site selection in Pampa La Bola, Chile last year, efforts have been substantially intensified in the host country of Chile to cultivate a strong relationship between the Observatory and local communities, with particular focus on engaging educators and students. In this way, the project not only broadens public understanding of high-energy astrophysics but also nurtures a sense of shared ownership and pride in scientific discovery. This engagement is a vital component of SWGO’s mission as it brings together cutting-edge research and the educational and cultural contexts of the region.
Our outreach efforts span a variety of educational initiatives that target students, teachers, and the broader public. In the short term, SWGO will provide training for teachers in physics and astronomy, with the objective of equipping local school educators with the tools and knowledge necessary to inspire the next generation of scientists. Members of the collaboration have already initiated contact with high school physics teachers from the nearby towns of San Pedro de Atacama and Toconao, to discuss ways to introduce concepts in astroparticle physics and the SWGO project itself. Recently, with support from the Atacama Astronomical Park (AAP), local teachers attended a summer school at Universidad Metropolitana de Ciencias de la Educación (UMCE) focused on the Sun–Earth–Moon system and its role in the Chilean elementary school curriculum. In the mid-term, we will implement a curriculum-linked science program for elementary and high school students in local towns, providing engaging hands-on experiences that make scientific concepts accessible and exciting. Keeping in line with our commitment to foster scientific curiosity in young people, we also plan to offer technical and scientific training, as well as financial support to advanced high school students. This initiative aims to ignite interest in pursuing STEM careers, particularly in fields such as physics and engineering.
Attempts are ongoing to translate and publish key content from the SWGO website in to Spanish, with the aim of making the project more accessible to the public in the host country.
For the broader community, the SWGO Scientific Literacy Program will offer an open-access course introducing key scientific concepts such as the scientific method, particle physics, and gamma rays. These efforts will demystify complex topics, address common misconceptions, and improve public understanding of the universe and the science that explores it.
Finally, SWGO is exploring the potential creation of a Community Learning Center, which would serve as an immersive educational hub for scientific tourism, possibly in collaboration with other observatories in the area. This center would offer interactive exhibits and activities, such as hands-on experiments and live demonstrations, to attract visitors eager to learn about the groundbreaking research at the observatory. This dynamic environment would also involve SWGO collaborators, who would share their work and its impact through presentations, videos, and educational materials. While still under consideration, the development of this center represents SWGO’s commitment to increasing community engagement and promoting the importance of science in daily life.
11 Summary
The SWGO project is now moving towards construction and will become the second major facility for ground-based gamma-ray astronomy in southern hemisphere, besides the CTAO Southern Array. Here we have shown the broad scientific impact to be expected from SWGO.
Its initial phase, SWGO-A, will already address key science questions through observations of the Galactic Center, bright Galactic sources including young remnants such as RX J1713.73946, SN 1006, and massive stellar clusters such as Westerlund 1. Source detections by SWGO-A within its first year of operation are expected to exceed the size of 3HWC and 1LHAASO catalogs. It can also provide particle physics constraints comparable to those from HAWC and LHAASO, but from a complementary region of the sky.
Highlights of the science case for SWGO include tightly constraining the thermal relic WIMP paradigm for dark matter, an expected rate of GRB detection of 0.7 per year, and probing Galactic source populations with PeV reach. In all of these cases the combination of SWGO and CTAO data can greatly enhance the scientific capabilities beyond what can be reached by either observatory alone. As a large array with unprecedented muon tagging capability, SWGO will have not just excellent background rejection power, but also the capability to make the first mass-resolved cosmic ray anisotropy measurements in this energy range and help to solve the muon puzzle in UHE air showers.
As we finalize the design of SWGO and prepare for construction, we are seeking partners from across the MM/MWL community, to further develop the SWGO science case and develop plans for operations.
Acknowledgments
The SWGO Collaboration gratefully acknowledges the support from a number of grants, agencies, and organizations that can be found listed here:
https://www.swgo.org/SWGOWiki/doku.php?id=acknowledgements.